### v1

# Investigating Emotions, Social Roles, Well-being and Mobile Usage Behaviors During COVID-19 Home Confinement

Abstract.
Pandemics and epidemics have been an essential part of human history. It has powerfully shaped human civilization, affecting politics, economics, medicine, and everyday human lives. Home confinement, one of the social isolation restrictions, has proven to be an effective way to prevent a pandemic and significantly reduce the spread of the virus. However, these measures have also led to negative psychological effects, which can be more severe than the physical discomfort caused by the pandemic itself. To address this issue, it is crucial to understand how being confined at home affects people’s psychological states and their use of digital devices. To this end, we conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks. We examined human emotions, social roles, well-being, and mobile usage behaviors and explored the predictability of affective states using digital traces. Extensive experimental results show that COVID-19 home confinement is associated with decreased valence and increased work time, and that human well-being could be accurately inferred from mobile usage behaviors. We also present a set of interesting findings on how different factors of mobile usage affect human emotions. Overall, the findings of this research have great potential to inform the development of interventions and remote programs to support individuals in similar situations, such as those who are ill or undergoing post-operative home rehabilitation and home work.
Emotion, COVID-19 Lockdown, Home Confinement, Mobile Sensing, Social Roles, Well-being
copyright: acmcopyrightjournalyear: 2023doi: 10.1145/1122445.1122456journal: IMWUTjournalvolume: 37journalnumber: 4article: 52publicationmonth: 8ccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
1. Introduction
The coronavirus disease 2019 (COVID-19) pandemic has resulted in a global lockdown, with many people experiencing home confinement for extended periods. Home confinement, defined as the restriction of individuals’ movements and activities within their homes (59), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (43; 16). Previous research has shown that home confinement may have negative psychological effects, including stress, anxiety, and depression. (6; 73). For instance, Bartoszek et al. (6) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue. Understanding the impact of home confinement on mental well-being is essential for developing interventions to support individuals during prolonged periods of confinement.
Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (73; 28). However, these approaches can be time-consuming and labor-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (29; 66)). Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being (68; 20). For instance, Rashid et al. (68) modeled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al. (20) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices. However, most sensing technologies are designed to monitor daily lives in general, rather than specifically for home confinement situations. As such, the patterns and daily behaviors captured by these technologies may differ significantly from those experienced during home confinement.
Studying the in-situ mental well-being and mobile usage behaviors of individuals confined to their homes can be challenging due to the variability of their degree of confinement, which may affect the accuracy and validity of collected data. Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement. On one hand, the strict measures put in place have created similar confinement conditions for all participants due to policy mandates. On the other hand, the COVID-19 home confinement has led to an increasing reliance of digital technologies(64), with srategies such as stay-at-home and work-from-home being recommended worldwide (69). As a result, people are increasingly turning to smartphones for social networks, online meetings, Internet surfing, or video games (69). Recent research shows that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (77). The reliance on technology during the COVID-19 lockdown has provided an excellent way to understand people’s mental well-being during home confinement situations.
In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during COVID-19 home confinement. Specifically, we are focusing on emotions and social roles because these aspects have a significant impact on overall well-being (22; 61). To assess emotional states, we utilize the widely accepted valence and arousal dimensions within the Circumplex Model of Affect (74) to assess emotional states. Social roles, as defined by Role Theory (8), are associated with norms, expectations, rules and behaviors that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance. This, in turn, can have a significant impact on an individual’s overall well-being and mental health. Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behavior. As a result, our research questions are:
  • RQ1. What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?
  • RQ2. Is it possible to unsupervisedly extract user activity from app usage records? If so, how are these activities related to emotions, productivity and interruptions?
  • RQ3. Can we model people’s emotion from mobile usage during home confinement? If so, what are the most significant types of features for prediction?
To answer the research questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia. To answer RQ1, we analyzed patterns of perceived valence, arousal and social roles during COVID-19 home confinement using Experience Sampling Method (ESM) (7). We also investigated the influence of the COVID-19 lockdown on emotions, social roles, and work-life balance by examining the end-of-day survey data. To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day. Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviors for RQ3. All scripts for this study are open source. 111link to be shared after the double-blind review. In summary, our contributions are as below:
  • We collected a comprehensive dataset from 32 participants over a period of 3 weeks during COVID-19 home confinement in Melbourne, Australia. The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc. Their smartphone and desktop usage were recorded through an app running in the background. The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
  • We conducted an in-depth analysis of the participants’ perceived valence, arousal, and social roles during the COVID-19 home confinement, and explore how lockdown affects people’s well-being and work-life balance. We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
  • We proposed an unsupervised activity recognition method called App-based Information Gain Temporal Segmentation (AIGTS). This method enables the extraction of user activities from app usage records. By identifying different segments of app sequences and applying clustering techniques, we were able to discover 16 distinct user activities in an unsupervised manner. We also discussed the correlations between these user activities and human well-being.
  • We conducted extensive experiments to predict human well-being based on smartphone usage behaviors. Experimental results demonstrated the superior prediction performance of the proposed model on mental well-being. Additionally, we derived different factors that contribute to understanding well-being and discussed the top predictors of human emotion during home confinement.
The remainder of the paper is as below. Section  2 presents related work on the impact of COVID-19 home confinement on human well-being, as well as recent advances in mobile sensing technologies for predicting well-being. Section  3 describes the data collection procedures. In Section 4, we explore patterns of human well-being during COVID-19 home confinement and investigate how the lockdown affects emotion and work-life balance. Next, we extract user activities through app sequences and explore correlations between user activities and human well-being in Section 5. Then, in Section  6, we present experimental results for predicting mental well-being during home confinement and provide a detailed discussion of the findings. Section 7 discusses the implications and Section 8 lists the limitations of our work. Finally, we summarize this research in Section  9 and indicate the potential direction of future work.
2. Related work
2.1. Impact of the Pandemic Restrictions on Human Well-being
Pandemics and epidemics have played an essential role throughout human history (82). In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world. Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (82; 19; 11), including post-traumatic stress symptoms (71), depression (32), stress (25), emotional disturbance (92), irritability(44), low mood (44), confusion (15), anger (52) and anxiety-induced insomnia (25). Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (38).
According to the World Health Organization  (89), the COVID-19 pandemic has severally impacted human well-being. The potential for physically sick or discomfort from COVID-19 is considered less of concern than the psychological and social impact (35). A growing body of literature has explored the impact of the COVID-19 on emotional and mental health (90; 21; 45; 42). Daly et al.  (21) explored the effects of COVID-19 on mental health in the UK. They found a significant increase in observed mental health problems, particularly for those aged 18-34 and women. Fountoulakis et al. (26) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased in 10.40%. Khubchandani et al. (41) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after initial COVID-19 lockdown in the USA. The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019. Killgore et al. (42) found that loneliness has increased during the COVID-19 pandemic, especially for those under lockdown restrictions.
Despite the existing research on the impact of pandemic restrictions on human well-being, much of it is limited to specific regions and conducted during different phases of varying degrees of restrictive measures. This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being. In light of this, the first research question (RQ1) focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage 4) were implemented.
2.2. Impact of the Pandemic Restrictions on Smartphone Usage
The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (84). Twitter even made an announcement to make working remotely a permanent option (12). As people perform their duties by staying at home or working from home, they increasingly turned to digital devices to communicate, interact and conduct job responsibilities (69). Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (77; 64). A recent global study (77) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (69).
Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (69; 13; 4). Ratan et al. (69) pointed out that problematic smartphone usage affects the ‘reward system’ of the brain and subsequent harms will persist in the form of various emotional disorders even after the lockdown effects are lifted. Arora et al. (4) explored the problematic phone use and its impact on mental health during the COVID-19 lockdown, and found that ‘Sadness’ and ‘anger’ were the dominant emotions in addictive Twitter users, while ‘valence’ was the dominating emotion among non-addictive users. Busch et al. (13) revealed that excessive mobile phone use may severely impact people’s mental health and human well-being. Especially during the COVID-19 lockdown, people’s mobile phone usage differs than normal, which increases the risk of overusing smartphones and affects their mental health (5).
2.3. Investigating User Activities from App Sequences
By analysing app sequences, researchers have been able to gain insights into various aspects of individuals’ behaviours in daily lives, including social roles  (2) and interruptiblity (62; 63). Deploying the approach ‘Application as a Sensor’, Anderson et al.  (2) found that app sequences on smartphones are correlated with individuals’ private and work social roles. Similarly, Okoshi et al. (62; 63) explored app sequences data and applied machine learning techniques to predict when a break-point in activity occurs for intelligent notification delivery. These approaches focus on identifying breakpoints between apps as a means to understand interruptibility. However, none of the previous research studies have utilised breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period. However, those approaches fail to capture the nuances and transitions between different app activities, limiting understanding of user behaviours. Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data. Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information pertaining to user behaviors and mental well-being. With this in mind, our research question RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
2.4. Inferring Human Well-being Using Mobile Sensing Technologies
The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (51) and psychological stages, e.g., valence (37), mood (56; 93; 55), engagement (30; 36), social anxiety (20; 68), depression (86; 83), etc. Most studies integrate data from various sources for building comprehensive prediction models. For instance, Gao et al. (30) predicted emotional, cognitive and behavioral student learning engagement in an Australian high school using wearable sensors and environmental sensors. Wang et al. (86) measured depression of undergraduate students at Dartmouth College using passive mobile sensing data and wearable data. Jaques et al. (37) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behavior data collected from mobile devices. Morshed et al. (56) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behavior, e.g., physical activity, communication patterns, and location. For example, Hintze et al. (34) analyzed mobile device usage characteristics such as session length, interaction frequency and daily usage in locked/unlocked state with respect to location context and diurnal pattern, to gain insights into how people use their smartphones and tablets. Madan (51) et al. designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (68; 83; 20), mood (93; 55), emotion (81), social roles (2) and interruptibility (62; 63). Zhang et al. (93) proposed an Android app to detect compound emotion utilizing smartphone usage data. Meegahapola et al. (55) explored the effect of geographical diversity on mood inference. They examined eight countries and compared the continent-specific, country-agnostic, and multi-country approaches trained for two mood inference tasks. Tlachac et al. (83) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine learning models to predict anxiety and depression. Tag et al. (81) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings. Rashid et al. (68) utilized sparsely collected mobile sensor data to predict perceived social anxiety. Chowed et al. (20) explored the fine-grained GPS data collected through mobile devices to model depression, social anxiety, affect, and social isolation.
While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement. Unlike daily life situations, individuals’ behaviors may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working. Although a few studies have examined psychological states using mobile sensing during the COVID-19 pandemic (83; 60), these works largely focus on broader environments rather than strictly home confinement. Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, aging or disability-related limited mobility. Therefore, we aim to address RQ3 by exploring the possibility of modeling emotions during home confinement and identifying significant features for prediction.
3. Data Collection
We conducted a field study in Melbourne, Australia, for three weeks, from 24 August 2020 to 13 September 2020. The study has been approved by the Human Research Ethics Committee at the authors’ university. Prior to the data collection, we obtained informed consent from all participants, ensuring that they were aware of the study’s purpose and that the dataset would be publicly released. The details about the background, participants, and collected data will be described in this section.
3.1. Background
Figure 1. The periods of six lockdowns and our data collection
To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (54; 80). Our data collection was conducted during the second lockdown period with Stage 4 restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage 4 restrictions in metropolitan Melbourne (58).
Under the Stage 4 restrictions in force at the time our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (69; 77). This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviors of individuals during home confinement. The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
(a) Age and gender
(b) Smartphone and desktop
Figure 2. Demographic and device usage information of participants
3.2. Participants
Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers. In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, std = 8.95). The age distribution of participants across genders can be found in Figure 2. It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample. While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behavior. Rather, we focus on individuals’ psychological states and smartphone usage behaviors during COVID-19 home confinement. Therefore, we believe that the potential oversample threat to the validity of our results is minimized. However, it is still important to note that our findings are specific to this particular population subset and may not be necessarily generalizable to other populations. Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
After the data collection, each participant was compensated a 25 AUD gift card as a token of appreciation for their time. Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage. Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants. Analysis of desktop usage behaviors will be included at the end of the discussion and explored further in future work.
3.3. Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (33) on their smartphones and desktops, respectively. These applications were specifically designed to record usage behaviors and enable background sensing and experience sampling, drawing inspiration from Van et al. (85). Both applications were optimized to minimize their consumption of computational power, energy, and storage resources. The applications utilized ESM, popping up a short questionnaire to participants every 90 minutes when their devices were actively in use between 7 am and 10 pm. This helps to obtain unbiased samples of participants’ activities, which served as the ground truth for our predictive models. Furthermore, an event-based ESM was triggered on the user’s smartphone after a minimum interaction of 10 minutes, with subsequent event-based ESMs occurring every 30 minutes.
At the end of each day, participants received an end-of-day survey via email at 7:30 pm. This survey aimed to gather reflections on their well-being and the impact of the lockdown on their daily experiences. This provides valuable context and additional data beyond the ESM questionnaires. Overall, we collected 502,485 recordings for smartphone usage, 1,749 ESM responses, and 265 end-of-day responses.
Table 1. An overview of the collected data
Category Items Questions / Descriptions Options
Smartphone Usage Application usage Foreground application name, package identifier, category N/A
Notifications Notification arrival time, content (hashed), sender (hashed), relationship N/A
Phone’s states Screen status, battery, ringer modes, GPS location, physical activities N/A
ESM Valence How did you feel in the last hour? (Unhappy vs Happy) 5-Likert scale
Arousal How did you feel in the last hour? (Calm vs Excited) 5-Likert scale
Social role In the last 15 minutes, I was engaged in: Work/private/both
Interruptibility I am currently available for other matters: Work/private/both/none
Task The task I managed to progress the most in the last 90 minutes: Multiple choices
End-of-day Survey Productivity I feel my day was productive. 5-Likert scale
Interruptibility I was continually interrupted today. 5-Likert scale
Valence Has the lockdown influenced your happiness today? 5-Likert scale
Arousal Has the lockdown influenced your energy today? 5-Likert scale
Work-life balance Has the lockdown impacted your work-life balance today? 5-Likert scale
Valence How has the lockdown influenced your happiness today? Text (optional)
Arousal How has the lockdown influenced your energy today? Text (optional)
Work-life balance How has the lockdown influenced your work life balance today? Text (optional)
Difference How different do you feel today was from yesterday? 5-Likert scale
Difference How was today different from yesterday? Text (optional)
3.3.1. Smartphone usage data
During the data collection, the background services of smartphones and desktops continuously tracked the following information: application usage (using Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (using Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (i.e. screen/battery status, ringer modes, location updates and physical activities). Specifically, the mobile app usage behavior was recorded every time the user opens the app. For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none). Table 1 displays an overview of the collected data for smartphone usage, ESM and end-of-day survey. In total, we have collected 502,485 recordings for smartphone usage data from 27 participants.
3.3.2. ESM responses
There are some popular questionnaires to measure well-being (39; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users. Especially, the item ‘valence’ was rated with a 5-point Likert scale from ‘unhappy’ to ‘happy’, and the item ‘arousal’ was also rated with a 5-point Likert scale from ‘calm’ to ‘excited’. Both valence and arousal were used to represent the users’ emotions based on the circumplex model (74). Understanding the social role, interruptiblity and task help us understand how people manage their time and attention on the digital devices (3). The detailed items of the ESM can be found in Table 1 and the screenshot of the ESM on users’ smartphones was shown in Figure 3(a). Overall, we have collected 1,749 ESM responses from 29 participants. Of these, 1,522 responses were obtained from 25 participants using smartphones (2 participants did not contribute any ESM responses), while the remaining 227 responses were from 10 participants using desktops.
(a) ESM
(b) End-of-day survey
Figure 3. The screenshots for the ESM and End-of-day survey
3.3.3. End-of-day survey
As shown in Figure 3(b), the end-of-day surveys were optional and included ten items to understand participants’ well-being during the lockdown such as productivity, valence, energy, and work-life balance. In addition, we surveyed how participants feel today compared to yesterday. This information could describe the dynamics of people’s experience during the COVID-19 lockdown. In the end-of-day survey, each item was rated with a 5-Likert scale or answered with free text input (see Table  1). Specifically, item 1 and item 2 were rated from 1 to 5 which indicates ‘strongly agree’, ‘somewhat agree’, ‘neither agree nor disagree’, ‘somewhat disagree’, ‘strongly disagree’, and items 3, 4, 5, 9 were rated from 1 to 5 which indicates ‘a great deal’, ‘a lot’, ‘a moderate amount’, ‘a little’, ‘not at all’. In total, we collected 265 responses from 25 participants during the data collection period.
4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
In this section, we aim to answer RQ1: ”What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”. To achieve this, we analyze 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1. Furthermore, we explore the influence of COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
4.1. The Perceived Valence, Arousal and Social roles
The distribution of perceived valence, arousal, and social roles from the ESM responses is depicted in Figure 4. On average, participants reported a valence level of 3.46 (SD = 1.01) and an arousal level of 2.72 (SD = 1.11). Usually, the relatively high valance and low arousal refer to a state of calmness, relaxation and peace. This could be attributed to various factors during the COVID-19 home confinement. One possible reason is that people may feel relieved to be safe at home, and cherish the opportunity to spend more quality time with their family and loved ones, away from the potential risks of infection outside. Furthermore, the absence of the usual stressors of daily life, such as commuting, work pressures, and social obligations, may be contributing to a general sense of calm and well-being. Additionally, some people may have found comfort in the sense of community and solidarity that emerged during the pandemic, as people came together to support one another through difficult times. All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement. Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage 4 restrictions (strict home confinement) were lifted and steps for reopening commenced. Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges. For instance, prolonged periods of confinement may lead to feelings of boredom, loneliness, and frustration.
(a) Valence and arousal
(b) Social roles
(c) Social roles and emotion
Figure 4. The distribution of perceived valence, arousal, social roles and their relationship
Next, we investigate the distribution of social roles and their correlation with emotion during home confinement. A social role refers to a set of norms, expectations, rules, beliefs and behaviors that are associated with a particular social situation (70). These roles can significantly influence users’ behaviors and the type of information they prefer to receive (3). As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
Since social roles play a crucial role in shaping individuals’ well-being, we investigate the emotions of users in different social roles, as shown in Figure 4(c). We found that, when engaging in private-related roles, participants had the highest levels of valence and lowest levels of arousal compared to work-related roles and hybrid roles. This may be because private-related roles involve personal activities which can provide a sense of comfort and happiness, while work-related roles or hybrid roles may involve stress, deadlines, and pressure to perform. The results of Welch’s t-tests (94) further support these findings, indicating that there is a significant difference in valence levels between private-related roles and the other two roles (i.e., private vs. both, private vs. work), with p-value less than 0.05. The above findings provides insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (31).
4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
In total, 54.3%, 51.3% and 54.7% of end-of-day responses displayed that the lockdown has influenced their valence, energy, and work-life balance to some extent, as shown in Figure 5. For the influence of lockdown on valence, 13 participants reported that it had a negative effect, claiming they ‘felt quite anxious and sad’ as ‘staying at home has made me a little sad’. They complained that they were unable to see friends (9 responses), see families (5 responses), go outside (10 responses), and feel bored (6 responses), e.g., ‘I thought I might be able to see friends, but I did not’, ‘I want to go outside, to the park’, ‘I’m missing my family’, ‘It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped ’. Nevertheless, two felt the lockdown has a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
(a) Productivity
(b) Interrupted
(c) Influence on well-being
Figure 5. The distribution of productivity, interruption, and influence during the lockdown
Regarding the impact of lockdown on arousal, 11 participants claimed that the impact was negative, reporting feeling tired (11 responses), being lazy (5 responses) and losing motivation (3 responses), e.g., ‘I don’t feel as motivated to do productive things’, ‘Emotionally tired, perhaps as a result of the lockdown’. However, two participants claimed the lockdown gave them ‘more energy’. For the influence on work-life-balance, most participants reported that they faced more interruptions or distractions (14 responses), had difficulty separating work and life (11 responses), spent more time on work due to interruptions or blurring of personal and work boundaries (15 responses), e.g., ‘I worked a bit more than I should’ve’, ‘Everything blends together, separating work and life is really difficult’, ‘Continuous interruptions, even my cat interrupts me’, ‘Need to work longer hours due to interruptions’. Overall, while a minority of responses showed that lockdown has a positive impact on their lives (e.g., more time with family or gardening), most responses showed negative influences of lockdown on their valence, arousal and work-life balance.
Besides the qualitative feedback from participants, we adopted Spearman correlation (23) to study the relationship between the reported Influence on Valence (from end-of-day survey) and the average perceived valence for that day (from ESM responses). In this way, we can determine whether the lockdown had a positive or negative effect on valence. Likewise, we went through the same process in terms of users’ arousal and social roles. For social roles, we correlate the reported Influence on wlb (‘wlf’ indicates work-life balance) to the percentage of the day spent on working (reported in ‘work’ or ‘both’ states). For example, if someone claimed that the lockdown significantly influenced their work-life balance (i.e., a score of 5, which becomes 3 after Likert transformation in the pre-processing), and they spent 50% of the day at work, then (50%, 3) would be a point in the correlation.
5. Understanding App Usage Behaviors During COVID-19 Home confinement
In this section, we address RQ2 ”Is it possible to unsupervisedly extract user activity from app usage records? If so, how are these activities related to emotions, productivity and interruptions?”. First, we introduce the process of app categorization. Next, we conduct unsupervised activity recognition from app sequences and display the average activities for all clusters. In this end, we explore the correlation between user activities and perceived psychological metrics.
Table 2. App categories and top three frequently used apps
Upper Category Category Top App Second App Third App
Business Business Teams LinkedIn Zoom
Communication Communication Chrome Discord WeChat
Entertainment Action Among Us Among Us
Casual Scrap Collector IdleCourier
Comics TachiyomiJ2K
Travel & Local Maps PTV
Video Players & Editors YouTube Bilibili YouTube Vanced
Entertainment Wow Steam RainbowSix
Food & Drink Uber Eats Menulog Mymacca’s
Health & Fitness Samsung Health Nike Training
Photography Photos Gallery Photos
Music & Audio Google Play Music Spotify YouTube Music
Role Playing Fate/GO
Shopping BIGAU Gumtree Depop
Productivity Productivity Excel Outlook PowerPoint
Education Canvas Student
Reading Books & Reference Audible AniDroid MendeleyDesktop
News & Magazines Joey Twitter Sync Dev
Social Social Facebook Instagram Snapchat
Tools Tools Gboard Explorer Samsung Keyboard
Others Auto & Vehicles Android Auto
Finance CommBank CommSec Up
House & Home Domain Realestate
Lifestyle Samsung Pay Tinder Hue
Medical MyTherapy E4 realtime
Parenting FamilyAlbum
Personalization One UI Home OnePlus Launcher TouchWiz home
5.1. App Categorization
We categorized the apps using historical usage logs. For Android apps, we referred to the categories listed in the Google Play Store, excluding uncategorized apps. Windows apps were manually sorted into the same categories as Google Play Store apps. Less frequently used apps, with a cumulative usage time of less than one hour across all participants, were excluded. Overall, we recorded 28 app categories and consolidated them into 8 upper-level categories: Business, Communication, Entertainment, Productivity, Reading, Social, Tools and Others. A summary of the top three most frequently used apps in each category is provided in Table 2. Notably, Chrome is categorized as a Communication app due to the scarcity of browser apps in the Google Play Store.
Categorizing desktop apps within the same categories used for mobile apps proves challenging, as they are often used for work and software development (e.g., Pycharm, WinSCP etc). In light of this, we grouped several development and utility apps under the Tools Category. Overall, the top 10 Tools apps for both desktop and mobile include: Gboard, Explorer, Samsung Keyboard, Always on Display, WinSCP, Pixel Launcher, Acrobat, pychar64, Google and Samsung keyboard. It is worth mentioning that developer tools such as Pycharm and WinSCP, as well as the utility Explorer, fall under this category.
5.2. Unsupervised Activity Recognition From App Sequences
Segmentation. To study how the application usage is related to the aforementioned variables, we propose App-based Information Gain Temporal Segmentation (AIGTS), which focuses on differentiating activities within app usage. Figure 6 provides an illustration of how activities are extracted from app usage data. AIGTS draws inspiration from Information Gain Temporal Segmentation (IGTS) (75), an unsupervised method for segmenting multivariate time series data. IGTS works by identifying which splits in the multivariate time series minimize the weighted entropy across the series. The variables within the multivariate time series data corresponded to whether an app of a particular category was used.
Unlike IGTS focusing solely on the time series, AIGTS aims to identify different segments and transition times between apps throughout the day. It is capable of detecting when participants change their app usage patterns, such as transitioning from using Email and Browsers (both categorized under Communication) to browsing Reddit (News and Media category). This method allows for segmentation at any level of granularity desired, ranging from half a day to as little as half a minute for application activity. The granularity of segmentation is determined by the variable $k$, which represents the number of splits made in the time series. To ensure that segments are captured at a consistent level of granularity, we propose the following steps for calculating the optimal $k$ in AIGTS:
  1. (1)
    Calculate the knee point of the activity data for each day;
  2. (2)
    Identify the average number of segments per minute for all days based on the knee point;
  3. (3)
    Compute the median average segments per minute for all users;
  4. (4)
    Multiply this median by the number of minutes per day (which varies from the first log to the last log) and choose this value as $k$.
These procedures guarantee that the average length of each segment will be consistent across all days.
Figure 6. An illustration of unsupervised activity recognition using app usage data
Clustering. Next, AIGTS was conducted on all time series, resulting in a segmented list for each user. To extract features from these segments, k-means clustering (47) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette analysis (48) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
Figure 7. Average activities for all clusters
Distribution of User Activities. Following the segmentation and clustering, we uncovered 16 clusters. The average duration of these user activities was 18 minutes. Figure 7 displays average activities for all clusters, where the average activity represents the average distribution of app categories appearing in each segment within the cluster. These clusters were manually named to describe the distribution of average activities. The $x$-axis displays the app categories that frequently appear in clusters, with infrequently appearing categories omitted. Notably, the cluster Other includes all uncategorizable segments and has roughly equal means across multiple categories. Additionally, the cluster Idling is used for Personalization apps because Personalization apps are typically launchers (see table 2), and indicate that the user is sitting on their device’s home screen. Moreover, the most frequent user activities during the COVID-19 lockdown were Just Communication and Idling, accounting for 17% and 14.60% of activities, respectively. Conversely, the least frequent activities were Business Communication and Music, accounting for 1.22% and 1.65% of activities, respectively.
5.3. App Usage Correlation
Furthermore, we investigate the correlations between user activities with their valence, arousal, productivity, and interruptions. To understand whether any particular activity $A$ was correlated with valence, we ran a Spearman correlation (23) between valence scores and the amount of time spent in user activity $A$ in the previous hour. The same goes for the arousal, productivity, and interruptions, but the productivity and interruptions vary in duration from 1 hour to 24 hours, as these are calculated from the end-of-day survey and correspond to the value for a full day. Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day. The top correlation results for how different activities relate to usesr are displayed in Table 3. We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
As shown in Table 3, the correlations are weak with most activities, but few activities have moderate correlations. Tools related activities were positively correlated with overall valence and for both age groups. Particularly, Idle Tools have a relatively stronger correlation with valence in older adults, indicating an increase in the mood at work (Idle Tools include phone home screens and phone keyboards). Surfing by women and young age group are negatively correlated to valence, indicating reading news or surfing (e.g., reading COVID-19-related news) are negatively related to their valence. Productivity was negatively correlated with Communication Tools (e.g., keyboards and chat clients), and positively correlated with Business Communication (i.e., both Communication and Business apps such as Microsoft Teams). It can be seen that while both are Communication applications, whether or not they are used for work may imply productivity levels. Interruptions can be seen as the inverse of productivity, as it has positive correlation with Communication Tools, but the value is higher. The correlation reaches a moderate level for older adults while weakening for the young age group. This might cause by the difference in working loads between young and older age groups. Lifestyle negatively correlates with interruptions, especially for female and older adults.
Table 3. Top correlations for all variables
Groups Variable Lowest correlation $\rho$ Highest correlation $\rho$
All Valence Work and Write (-0.07) Idle Tools (0.11)
Arousal Business (-0.05) Social Tools (0.15)
Productivity Communication Tools (-0.24) Business Communication (0.18)
Interruptions Lifestyle (-0.15) Communication Tools (0.29)
Male Valence Just Communication (-0.08) Business Communication (0.10)
Arousal Idling (-0.12) Just Communication (0.11)
Productivity Communication Tools (-0.15) Business Communication (0.34)
Interruptions Video Watching (-0.23) Communication Tools (0.18)
Female Valence Surfing (-0.17) Other (0.13)
Arousal Work and Write (-0.17) Social Tools (0.21)
Productivity Communication Tools (-0.30) Lifestyle (0.16)
Interruptions Lifestyle (-0.33) Idling (0.41)
Age 18-24 Valence Surfing (-0.13) Just Tools (0.09)
Arousal Idle Tools (-0.17) Social Tools (0.19)
Productivity Communication Tools (-0.21) Work and Write (0.20)
Interruptions Social Tools (-0.10) Communication Tools (0.16)
Age $\geq$25 Valence Just Communication (-0.12) Idle Tools (0.23)
Arousal Work and Write (-0.10) Idle Tools (0.15)
Productivity Idle Tools (-0.25) Social (0.20)
Interruptions Lifestyle (-0.24) Communication Tools (0.42)
Table 4. The computed features from smartphone usage behaviors
Category Feature Description
Contextual Information weekday The day of the week (e.g., Monday, Tuesday)
time_period The time of the day (morning, afternoon, evening and midnight)
sin_weekday Sine Transformation of weekday
cos_weekday Cosine Transformation of weekday
sin_time Sine Transformation of time period
cos_time Cosine Transformation of time period
is_weekend Binary values describing whether is is weekend or not
Notification num_noti Number of notifications
num_noti_interacted Number of notifications interacted
noti_cat Categorical Transformation of the number of notifications
noti_intr_cat Categorical Transformation of the number of notifications interacted
noti_interacted_pcg The percentage of notification interacted to the total number of notifications
num_communication Number of notifications from the Communication Application
main_contact The main social relationship that contacted with
Physical Activity nunique_activity Number of unique physical activities
main_activity The main physical activity of the users
num_activity Number of physical activities
App nunique_app The number of unique apps
num_app The total number of apps
nunique_app_cat The number of unique app categories
app_noti The number of notification generated by non-communication apps
nunique_app_noti The number of unique apps sent notifications
main_cat_app The main category of apps
second_cat_app The second main category of apps
Activity Inferred from Apps nunique_user_activity The number of unique user activities
num_user_activity The number of user activities
main_user_activity The main user activity
sec_user_activity The second user activity
entropy_user_activity The entropy of different categories of user activity
6. Predicting User Emotion and Social Roles from Mobile Usage Data
In this section, we address RQ3 ”Can we model people’s emotion from mobile usage during home confinement? If so, what are the most significant types of features for prediction?”. First, we introduce the features extracted from mobile usage data and the prediction pipeline. We then list and discuss the prediction results of the general models and individual models. Finally, we examine the significant features for predicting human well-being.
6.1. Feature Extraction
Table 4 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (46; 76; 91), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transforme them into sine and cosine values. Additionally, it is worth noting that app-usage patterns exhibit variations between weekdays and weekends (40; 46), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notification is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app-usages, we employed our proposed method, AIGTS, as detailed in Section 5.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combing two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combing the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.
6.2. Prediction Pipeline
while emotion prediction can be addresed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (65; 93). We believe that utilizing regression is a better approach becuase valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
Emotion. The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study. Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’. Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicates ‘calm’ and 5 means ‘excited’. To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
Regressors. In this research, we aim to develop regression models that accurately predicts human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (79) regressor , Random Forest (RF) (78) regressor, and Gradient Boosting (GB) (27) regressor. KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbors to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way thatminimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
Feature Selection. Various methods, such as SHAP (50), LIME (72), and the $F$-statistic ($F$) (67), can be used for feature selection. In this study, we employ the $F$-statistic to determine the appropriate number of features. While most of our features are numeric, the $F$-statistic measures the univariate linear dependence between each feature and the target variable, taking into account statistical significance. Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable. In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and features selected based on an $F$ value greater than or equal to 10.
Validation. To compensate for the limited richness of our data, we have implemented nested cross-validation (CV) (57) to train and evaluate our prediction model. This approach allows us to accurately estimate the unbiased generalization performance of the model. The nested CV consists of an inner loop and an outer loop. Initially, the dataset is divided into training, validation, and testing sets. This helps optimize the model’s performance. In the outer loop, similar to previous human-centered studies (30; 24), we divide the data into multiple groups, each group representing one participant. To ensure unbiased evaluation, we apply k-fold cross validation (88) where k=5 on all groups. This means that the data from the same participants are not used in both the training and test sets simultaneously. In the inner loop, the remaining data groups are further split into three folds, with each fold serving as a validation set. We optimize the hyperparameters on the training set, evaluate them on the validation set, and select the best parameter settings based on the results from the three folds.
(a) MAE for valence
(b) MAE for arousal
(c) RMSE for valence
(d) MAE for arousal
Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents a 95% confidence interval
Baselines. Similar to previous human-centered studies (30; 87), we compare the proposed model with two baselines. The first baseline is the Mean baseline, which calculates the average score of user responses. The second baseline is the Median baseline, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (18). Generally, lower values of MAE or RMSE indicate better prediction performance.
6.3. Prediction Result
To address RQ3, we proposed several regression models in the experiment. First, a general prediction model was constructed by considering all participants in the study. Second, we built individual prediction models for each participant to account for the unique characteristics and patterns exhibited by different individuals. It is worth mentioning that only mobile phone data was utilized in this study. After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis.
6.3.1. General Model
The prediction performance for the general models is presented in Figure 8. The models used in the comparison include Mean baseline, Median baseline, KNN regressor, RF regressor, GB regressor. Among these models, the RF regressor demonstrated superior performance to both baselines in terms of MAE and RMSE for predicting valance and arousal. Interestingly, the RF regressors with different feature sets achieved comparable performance, with all feature sets yielding the best results. This can be attributed to the capability of RF to assign weights to features based on their importance, where incorporating more features leads to improved prediction performance. Furthermore, it is observed that the proposed model generally exhibited better performance in predicting valence compared to arousal, the difference in performance may be attributed to the complexity of arousal, which is influenced by various factors such as physiological arousal and cognitive processing. These factors may not be fully captured by mobile usage behaviours alone.
Table 5. Prediction performance of the general models with different feature sets
Prediction Result (RMSE/MAE)
Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
All features
Valence (29) 0.986/0.824 1.091/0.767 0.963/0.776 0.903/0.717 0.925/0.743
Arousal (29) 1.113/0.884 1.146/0.784 1.088/0.844 1.005/0.766 1.084/0.756
Features with $F$-statistic ¿= 5
Valence (17) 0.986/0.824 1.091/0.767 0.956/0.769 0.904/0.717 0.987/0.725
Arousal (14) 1.113/0.884 1.146/0.784 1.064/0.818 1.015/0.778 1.087/0.747
Features with $F$-statistic ¿= 10
Valence (13) 0.986/0.824 1.091/0.767 0.936/0.730 0.912/0.725 0.987/0.724
Arousal (12) 1.113/0.884 1.146/0.784 1.064/0.819 1.011/0.774 1.076/0.745
Table 5 displays the prediction results of the general models for valence and arousal. Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets. For the regression model using all features, the RMSE obtained by the RF algorithm was 0.903, surpassing the Median baseline by 17.23%. Furthermore, the MAE of the RF regressor using all feature sets was 0.717, outperforming the mean baseline by 12.99% and the median baseline by 6.52%. Similar results were observed for the models utilizing the top 17 features. Regarding the arousal dimension, our results indicate that the RF regressor utilizing the all features achieved the best performance in terms of RMSE, with a value of 1.005. This outperformed the Median baseline by 12.30%. These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
6.3.2. Individual Models
The individual models for both valence and arousal were trained using all 29 features. However, due to some participants having insufficient data recorded, the models were only trained on those with more than 50 data points. This resulted in a total of 13 participants and 1220 data points being used. The prediction results for these models, with the ID of participants indicated on the y-axis, can be seen in Figure 9. Overall, we can observe variations in the prediction performance of valence and arousal across different participants. For instance, participant ‘M11OAR’ consistently exhibits higher MAE and RMSE on both valence and arousal than the other participants, e.g., ‘H12DPB’ and ‘E02OKG’. This variability suggests that individual differences play a significant role in determining the accuracy of the predictions, highlighting the need to consider personalized factors when developing prediction models for emotion in home confinement situations.
Additionally, it is worth noting that while the proposed valence models exhibited better prediction performance than baseline models for certain participants (e.g., ‘M02MMP’, ‘L10FGB’), there were participants who did not achieve satisfactory results. The possible reasons may be two-fold: firstly, individual differences play a crucial role in determining valence, and it may not be straightforward to accurately predict valence solely based on mobile usage behaviors for some participants. This highlights the complexity and subjectivity of valence and the need for a more nuanced understanding of individual emotional experiences; secondly, insufficient data availability may have limited the performance of the predictive models for some participants. Having a larger and more diverse dataset would enable better training and improve the performance of the predictive models. These findings contribute to our understanding of personalized prediction models for well-being, which could ultimately lead to more tailored interventions and support systems to enhance individuals’ well-being during periods of home confinement.
(a) MAE for valence
(b) MAE for arousal
(c) RMSE for valence
(d) RMSE for arousal
Figure 9. Prediction performance of individual models using all features
6.4. Feature Importance
To interpret the key factors influencing the prediction of valence and arousal during home confinement, we examine the feature importance of the best-performing models, which in this case are the RF regressors. For brevity, we focus on the top 15 features for each model, as shown in Figure 10. The RF regressors utilize impurity-based feature importance, which measures how much including a particular feature reduces impurity across the decision tree nodes. A higher score indicates a more important feature, as it leads to a greater reduction in impurity. A higher score indicates a more important feature, as it leads to a greater reduction in impurity. As depicted in Figure 10, we observe that the most important features for both valence and arousal are similar, but with different weights. These features primarily include time-related features, app-category features, and notification features. These findings suggest that these variables play a significant role in accurately predicting participants’ valence and arousal levels during home confinement.
(a) Valence
(b) Arousal
Figure 10. Top 15 features based on impurity score with RF regressors
Both models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day significantly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less significance in the arousal model. Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels. These findings differ slightly from the observations reported in (40), which found that app-usage patterns varied greatly vary by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (17), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
Above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement. They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods. By considering the impact of app usage, notifications, and time-related factors, we can also design targeted strategies to enhance individuals’ emotional well-being and overall quality of life.
7. Implications
Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance.Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles. The findings from our study are significant in raising awareness about the importance of mental health not only during periods of home confinement but also in similar situations such as post-operative recovery, parental leave, aging, or limited mobility related to disability. These findings can also play a crucial role in the early detection and treatment of mental illnesses that may arise during these circumstances. Additionally, policymakers can utilize our research to develop appropriate strategies that minimize the negative effects of pandemic restrictive policies on mental health. To achieve this, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al. (11).
In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (49). Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviors to those during the pandemic (53), underscoring the importance of studying people’s behaviors in high technology-reliance contexts. Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions. Furthermore, our study contributes to understanding the mental health benefits and consequences of working from home or relying heavily on technology. For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviors of individuals under home confinement conditions, which has important implications for various fields. For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation. Understanding how individuals cope with home confinement and the role that technology plays in their lives can help healthcare professionals or researchers develop effective interventions and design effective remote support programs to improve the overall quality of care for patients in similar situations. Furthermore, the findings from this study may have broader implications for our understanding of human behavior and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
8. Limitations
We have identified three main limitations in this research. One significant limitation is the selection of time and place. Regarding time, we were unable to compare user well-being and digital device usage data before COVID-19 lockdown with the data during the lockdown due to the unpredictable nature of lockdown policies. However, we have mitigated this limitation by allowing users to report the impact of the lockdown on their well-being and work-life balance through self-reported questionnaires. These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4. In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage 4 restrictions was applied. While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (10; 14; 11). Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
The second limitation is the relatively short duration of the data collection period. We chose to collect data for only three weeks in order to minimize interruptions from ESM questionnaires and end-of-day surveys, with the goal of reducing the burden on participants during the challenging realities of COVID-19 home confinement. Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses. It is worth noting that the delivery of end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience. Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data. However, due to the limited three-week period of our data collection, it is important to acknowledge that if strict home confinement lasts longer, the emotional experiences of individuals may evolve as they adapt to new challenges.
Finally, the third limitation of this study is the number and representativeness of the participants. A total of 32 participants were recruited, with 27 using smartphones, 13 using desktops, and only 8 using both. Since the majority of participants use mobile phones for data collection, the majority of our background application data analysis focuses on mobile usage data. Additionally, participants were primarily recruited through Facebook and Discord advertisements, which may have led to a population that is skewed towards tech-savvy information workers. Additionally, there is a possibility of oversampling individuals who were ‘relieved to be safe at home and cherish the opportunity to spend more quality time with their family,’ as these individuals may have had more availability to participate in the study. It is important to acknowledge that the COVID-19 lockdown presented challenges in recruiting participants. However, despite these challenges, we were able to collect sufficient data from each participant to conduct our analysis.
9. Conclusion
In this study, we delved into how human emotions, social roles, and mobile usage behaviors were affected during the COVID-19 home confinement period. Our proposed models yielded promising results in predicting human emotions during home confinement. Moreover, we uncovered intriguing insights regarding the correlation between different factors and human well-being. To address our research objectives on home confinement, we formulated three research questions and key findings are summarized below.
In response to RQ1, we found that participants reported relatively high levels of valence but relatively low levels of arousal, suggesting that participants experienced a state of calmness and relaxation during COVID-19 home confinement. Additionally, our analysis suggested that the COVID-19 lockdown has a detrimental effect on valence and work-life balance, with participants experiencing longer work hours due to interruptions and a blurring of personal and work boundaries. Regarding RQ2, we segmented app usage data and identified the most frequent user activities ‘Just Communication’ and ‘Idling’, potentially attributed to increased virtual communication and available free time. We also found that tools-related activities positively correlated with valence, suggesting people used productivity apps effectively. Moreover, interruptions displayed a positive association with Communication Tools, possibly due to the social nature of these tools leading to more interruptions during work hours. Moving on to RQ3, we developed regression models to predict people’s valence and arousal based on their mobile usage during home confinement. These models exhibited superior performance compared to baseline models. Notably, our findings revealed that time-related features emerged as the primary predictors for both valence and arousal.
Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviors during home confinement is crucial. This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health. Moving forward, our research will expand to investigate users’ desktop usage behavior. By comparing the differences in usage behavior across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
10. Acknowledgments
We sincerely thank the anonymous reviewers.

### v2

@@
-Investigating Emotions, Social Roles, Well-being and Mobile Usage Behaviors During COVID-19 Home Confinement
+”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement
 Abstract.
-Pandemics and epidemics have been an essential part of human history.
-It has powerfully shaped human civilization, affecting politics, economics, medicine, and everyday human lives.
-Home confinement, one of the social isolation restrictions, has proven to be an effective way to prevent a pandemic and significantly reduce the spread of the virus.
-However, these measures have also led to negative psychological effects, which can be more severe than the physical discomfort caused by the pandemic itself.
-To address this issue, it is crucial to understand how being confined at home affects people’s psychological states and their use of digital devices.
-To this end, we conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks.
-We examined human emotions, social roles, well-being, and mobile usage behaviors and explored the predictability of affective states using digital traces.
-Extensive experimental results show that COVID-19 home confinement is associated with decreased valence and increased work time, and that human well-being could be accurately inferred from mobile usage behaviors.
-We also present a set of interesting findings on how different factors of mobile usage affect human emotions.
-Overall, the findings of this research have great potential to inform the development of interventions and remote programs to support individuals in similar situations, such as those who are ill or undergoing post-operative home rehabilitation and home work.
-Emotion, COVID-19 Lockdown, Home Confinement, Mobile Sensing, Social Roles, Well-being
-copyright: acmcopyrightjournalyear: 2023doi: 10.1145/1122445.1122456journal: IMWUTjournalvolume: 37journalnumber: 4article: 52publicationmonth: 8ccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
+Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life.
+While essential in various scenarios, confinement within the home environment can profoundly influence psychological well-being and digital device usage.
+In this study, we delve into these effects, utilising the COVID-19 lockdown as a special case study to draw insights extending to various homebound situations.
+We conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks and analysed their emotions, well-being, social roles, and mobile usage behaviours.
+We extracted user activity from app usage records in an unsupervised manner, and experimental results revealed that app usage behaviours are effective indicators of emotional well-being in confined environments.
+Our research has great potential for developing supportive strategies and remote programs, not only for people facing similar medical isolation situations, but also for individuals in long-term home confinement, such as those with chronic illnesses, recovering from surgery, or adapting to permanent remote work arrangements.
+Emotion, Home Confinement, Mobile Sensing, COVID-19 Lockdown, Social Roles, Well-being
+copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXjournal: TOCHIccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
 1. Introduction
-The coronavirus disease 2019 (COVID-19) pandemic has resulted in a global lockdown, with many people experiencing home confinement for extended periods.
-Home confinement, defined as the restriction of individuals’ movements and activities within their homes (59), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (43; 16).
-Previous research has shown that home confinement may have negative psychological effects, including stress, anxiety, and depression.
-(6; 73).
+Home confinement, defined as the restriction of individuals’ movements and activities within their homes (63), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (47; 17).
+Previous research indicated that home confinement may adversely affect mental well-being, leading to stress, anxiety, and depression (7; 78).
 For instance, Bartoszek et al.
 (7) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue.
-Understanding the impact of home confinement on mental well-being is essential for developing interventions to support individuals during prolonged periods of confinement.
-Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (73; 28).
-However, these approaches can be time-consuming and labor-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (29; 66)).
-Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being (68; 20).
+Understanding the impact of home confinement on mental well-being is essential for developing effective interventions to support individuals during prolonged periods of confinement.
+Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (78; 30).
+However, these approaches can be time-consuming and labour-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (32; 71)).
+Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being unobtrusively (73; 21).
 For instance, Rashid et al.
-(68) modeled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al.
+(73) modelled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al.
 (21) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices.
-However, most sensing technologies are designed to monitor daily lives in general, rather than specifically for home confinement situations.
-As such, the patterns and daily behaviors captured by these technologies may differ significantly from those experienced during home confinement.
-Studying the in-situ mental well-being and mobile usage behaviors of individuals confined to their homes can be challenging due to the variability of their degree of confinement, which may affect the accuracy and validity of collected data.
+In home confinement scenarios, traditional mobile sensing methods, heavily relying on sensors like accelerometers and GPS to track physical behaviours, encounter unique challenges.
+These methods, which infer mental well-being from patterns such as visited locations, steps taken, and stationary periods (73), become less effective when individuals are confined at home.
+During home confinement, individuals’ interaction with their smartphones changes significantly.
+Rather than carrying their devices with them, people tend to leave their phones in one place.
+This shift in usage patterns diminishes the effectiveness of conventional metrics such as step counts and travel patterns, which are less indicative of mental well-being in a stationary context.
+Consequently, there is a shift towards analyzing more nuanced data types, such as application usage patterns, to better understand well-being under home confinement.
+Therefore, in this research, we focus on these finer-grained user activities to uncover deeper insights into the subtle changes in mental well-being during home confinement.
+Another challenge in studying the in-situ mental well-being of individuals confined to their homes is establishing a natural yet controlled environment.
+Home confinement situations vary significantly among individuals, influenced by different reasons and individual differences, making effective modelling challenging.
 Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement.
-On one hand, the strict measures put in place have created similar confinement conditions for all participants due to policy mandates.
-On the other hand, the COVID-19 home confinement has led to an increasing reliance of digital technologies(64), with srategies such as stay-at-home and work-from-home being recommended worldwide (69).
-As a result, people are increasingly turning to smartphones for social networks, online meetings, Internet surfing, or video games (69).
-Recent research shows that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (77).
-The reliance on technology during the COVID-19 lockdown has provided an excellent way to understand people’s mental well-being during home confinement situations.
-In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during COVID-19 home confinement.
-Specifically, we are focusing on emotions and social roles because these aspects have a significant impact on overall well-being (22; 61).
+The strict measures in place have created similar confinement conditions for all participants due to policy mandates.
+Additionally, the COVID-19 home confinement has led to an increasing reliance on digital technologies (68), with strategies such as stay-at-home and work-from-home being recommended worldwide (74).
+Recent research indicates that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (82).
+This reliance on technology provides an excellent way to understand people’s mental well-being during home confinement.
+In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during home confinement.
+Specifically, we focus on emotions and social roles, as these aspects significantly impact overall well-being (23; 65).
 To assess emotional states, we utilize the widely accepted Valence and Arousal dimensions within the Circumplex Model of Affect (79) to assess emotional states.
-Social roles, as defined by Role Theory (8), are associated with norms, expectations, rules and behaviors that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance.
+Social roles, as defined by Role Theory (9), involve norms, expectations, rules and behaviours that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance.
 This, in turn, can have a significant impact on an individual’s overall well-being and mental health.
-Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behavior.
+Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behaviours.
 As a result, our research questions are:
@@
 To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day.
-Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviors for RQ3.
+Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviours for RQ3.
 All scripts for this study are open source.
@@
 In summary, our contributions are as below:
-  • We collected a comprehensive dataset from 32 participants over a period of 3 weeks during COVID-19 home confinement in Melbourne, Australia.
+  • We collected a comprehensive dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia.
 The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc.
@@
 The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
-  • We conducted an in-depth analysis of the participants’ perceived valence, arousal, and social roles during the COVID-19 home confinement, and explore how lockdown affects people’s well-being and work-life balance.
+  • We analyzed participants’ perceived valence, arousal, and social roles during home confinement, and explored how lockdown affected well-being and work-life balance.
 We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
-  • We proposed an unsupervised activity recognition method called App-based Information Gain Temporal Segmentation (AIGTS).
-This method enables the extraction of user activities from app usage records.
-By identifying different segments of app sequences and applying clustering techniques, we were able to discover 16 distinct user activities in an unsupervised manner.
-We also discussed the correlations between these user activities and human well-being.
-  • We conducted extensive experiments to predict human well-being based on smartphone usage behaviors.
-Experimental results demonstrated the superior prediction performance of the proposed model on mental well-being.
-Additionally, we derived different factors that contribute to understanding well-being and discussed the top predictors of human emotion during home confinement.
-The remainder of the paper is as below.
-Section  2 presents related work on the impact of COVID-19 home confinement on human well-being, as well as recent advances in mobile sensing technologies for predicting well-being.
-Section  3 describes the data collection procedures.
-In Section 4, we explore patterns of human well-being during COVID-19 home confinement and investigate how the lockdown affects emotion and work-life balance.
-Next, we extract user activities through app sequences and explore correlations between user activities and human well-being in Section 5.
-Then, in Section  6, we present experimental results for predicting mental well-being during home confinement and provide a detailed discussion of the findings.
-Section 7 discusses the implications and Section 8 lists the limitations of our work.
-Finally, we summarize this research in Section  9 and indicate the potential direction of future work.
+  • We proposed an unsupervised activity recognition method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records.
+This approach autonomously identified 16 distinct user activities in an unsupervised manner, and we then explored the correlations between these user activities and human well-being.
+  • We conducted extensive experiments to predict human well-being in confined environments from smartphone usage behaviours, demonstrating the superior performance of our model.
+Additionally, we identified key factors contributing to well-being and discussed the primary predictors of human emotion during home confinement.
 2. Related work
@@
 In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world.
-Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (82; 19; 11), including post-traumatic stress symptoms (71), depression (32), stress (25), emotional disturbance (92), irritability(44), low mood (44), confusion (15), anger (52) and anxiety-induced insomnia (25).
+Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (87; 20; 12), including post-traumatic stress symptoms (76), depression (35), stress (26), emotional disturbance (97), irritability(48), low mood (48), confusion (16), anger (56) and anxiety-induced insomnia (26).
 Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (42).
-According to the World Health Organization  (89), the COVID-19 pandemic has severally impacted human well-being.
-The potential for physically sick or discomfort from COVID-19 is considered less of concern than the psychological and social impact (35).
-A growing body of literature has explored the impact of the COVID-19 on emotional and mental health (90; 21; 45; 42).
+According to the World Health Organization  (94), the COVID-19 pandemic has severely impacted human well-being.
+The potential for physically sick or discomfort from COVID-19 is considered less of a concern than the psychological and social impact (38).
+A growing body of literature has explored the impact of COVID-19 on emotional and mental health (95; 22; 49; 46).
 Daly et al.
-(21) explored the effects of COVID-19 on mental health in the UK.
-They found a significant increase in observed mental health problems, particularly for those aged 18-34 and women.
+(22) explored the effects of COVID-19 on mental health in the UK and found a significant increase in observed mental health problems.
 Fountoulakis et al.
-(26) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased in 10.40%.
+(28) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased by 10.40%.
 Khubchandani et al.
-(41) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after initial COVID-19 lockdown in the USA.
+(45) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after the initial COVID-19 lockdown in the USA.
 The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019.
@@
 This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being.
-In light of this, the first research question (RQ1) focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage 4) were implemented.
+In light of this, RQ1 focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage IV) were implemented.
 2.2. Impact of the Pandemic Restrictions on Smartphone Usage
-The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (84).
+The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (89; 27; 4).
 Twitter even made an announcement to make working remotely a permanent option (13).
-As people perform their duties by staying at home or working from home, they increasingly turned to digital devices to communicate, interact and conduct job responsibilities (69).
-Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (77; 64).
+As people perform their duties by staying at home or working from home, they increasingly turn to digital devices to communicate, interact and conduct job responsibilities (74).
+Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (82; 68).
 A recent global study (82) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (74).
-Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (69; 13; 4).
+Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (74; 14; 5).
 Ratan et al.
@@
 2.3. Investigating User Activities from App Sequences
-By analysing app sequences, researchers have been able to gain insights into various aspects of individuals’ behaviours in daily lives, including social roles  (2) and interruptiblity (62; 63).
+By analysing app sequences, researchers have gained insights into various aspects of individuals’ daily behaviours, including social roles  (2) and interruptiblity (66; 67).
 Deploying the approach ‘Application as a Sensor’, Anderson et al.
@@
 Similarly, Okoshi et al.
-(62; 63) explored app sequences data and applied machine learning techniques to predict when a break-point in activity occurs for intelligent notification delivery.
-These approaches focus on identifying breakpoints between apps as a means to understand interruptibility.
-However, none of the previous research studies have utilised breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
+(66; 67) utilized app sequence data and machine learning techniques to predict breakpoints in activity for intelligent notification delivery.
+These studies focus on identifying breakpoints between apps to understand interruptibility.
+However, no previous research has used breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
 Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period.
@@
 Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data.
-Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information pertaining to user behaviors and mental well-being.
-With this in mind, our research question RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
+Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information about user behaviours and mental well-being.
+With this in mind, RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
 2.4. Inferring Human Well-being Using Mobile Sensing Technologies
-The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (51) and psychological stages, e.g., valence (37), mood (56; 93; 55), engagement (30; 36), social anxiety (20; 68), depression (86; 83), etc.
+The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (55) and psychological stages, e.g., valence (41), mood (60; 98; 59), engagement (33; 39; 31), social anxiety (21; 73), depression (91; 88), etc.
 Most studies integrate data from various sources for building comprehensive prediction models.
 For instance, Gao et al.
-(30) predicted emotional, cognitive and behavioral student learning engagement in an Australian high school using wearable sensors and environmental sensors.
+(33) predicted emotional, cognitive and behavioural student learning engagement in an Australian high school using wearable sensors and environmental sensors.
 Wang et al.
@@
 Jaques et al.
-(37) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behavior data collected from mobile devices.
+(41) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behaviour data collected from mobile devices.
 Morshed et al.
 (60) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
-While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behavior, e.g., physical activity, communication patterns, and location.
+While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behaviour, e.g., physical activity, communication patterns, and location.
 For example, Hintze et al.
@@
 designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
-Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (68; 83; 20), mood (93; 55), emotion (81), social roles (2) and interruptibility (62; 63).
+Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (73; 88; 21), mood (98; 59), emotion (86), social roles (2) and interruptibility (66; 67).
 Zhang et al.
-(93) proposed an Android app to detect compound emotion utilizing smartphone usage data.
-Meegahapola et al.
+(98) proposed an Android app to detect compound emotion utilizing smartphone usage data and Meegahapola et al.
 (59) explored the effect of geographical diversity on mood inference.
-They examined eight countries and compared the continent-specific, country-agnostic, and multi-country approaches trained for two mood inference tasks.
 Tlachac et al.
-(83) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine learning models to predict anxiety and depression.
+(88) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine-learning models to predict anxiety and depression.
 Tag et al.
 (86) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings.
-Rashid et al.
-(68) utilized sparsely collected mobile sensor data to predict perceived social anxiety.
-Chowed et al.
-(20) explored the fine-grained GPS data collected through mobile devices to model depression, social anxiety, affect, and social isolation.
 While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement.
-Unlike daily life situations, individuals’ behaviors may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working.
-Although a few studies have examined psychological states using mobile sensing during the COVID-19 pandemic (83; 60), these works largely focus on broader environments rather than strictly home confinement.
-Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, aging or disability-related limited mobility.
-Therefore, we aim to address RQ3 by exploring the possibility of modeling emotions during home confinement and identifying significant features for prediction.
+Unlike daily life situations, individuals’ behaviours may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working.
+Although a few studies (88; 64) have examined psychological states using mobile sensing during the COVID-19 pandemic, these works largely focus on broader environments rather than strictly home confinement.
+Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, ageing or disability-related limited mobility.
+Therefore, we aim to address RQ3 by exploring the possibility of modelling emotions during home confinement and identifying significant features for prediction.
 3. Data Collection
@@
 Figure 1. The periods of six lockdowns and our data collection
-To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (54; 80).
-Our data collection was conducted during the second lockdown period with Stage 4 restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1).
-Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage 4 restrictions in metropolitan Melbourne (58).
-Under the Stage 4 restrictions in force at the time our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons.
+To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (58; 85).
+Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1).
+Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (62).
+Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons.
 Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning.
@@
 On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions.
-During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (69; 77).
-This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviors of individuals during home confinement.
+During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (74; 82).
+This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviours of individuals during home confinement.
 The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
@@
 It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample.
-While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behavior.
-Rather, we focus on individuals’ psychological states and smartphone usage behaviors during COVID-19 home confinement.
+While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behaviour.
+Rather, we focus on individuals’ psychological states and smartphone usage behaviours during COVID-19 home confinement.
 Therefore, we believe that the potential oversample threat to the validity of our results is minimized.
@@
 Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
-After the data collection, each participant was compensated a 25 AUD gift card as a token of appreciation for their time.
+After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time.
 Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage.
 Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants.
-Analysis of desktop usage behaviors will be included at the end of the discussion and explored further in future work.
+Analysis of desktop usage behaviours will be explored further in future work.
 3.3. Procedures
 Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (36) on their smartphones and desktops, respectively.
-These applications were specifically designed to record usage behaviors and enable background sensing and experience sampling, drawing inspiration from Van et al.
+These applications were specifically designed to record usage behaviours and enable background sensing and experience sampling, drawing inspiration from Van et al.
 (90).
@@
 screen/battery status, ringer modes, location updates and physical activities).
-Specifically, the mobile app usage behavior was recorded every time the user opens the app.
+Specifically, the mobile app usage behaviour was recorded every time the user opened the app.
 For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none).
@@
 3.3.2. ESM responses
-There are some popular questionnaires to measure well-being (39; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research.
+There are some popular questionnaires to measure well-being (43; 10), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (10) was employed in this research.
 In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users.
@@
 4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
-In this section, we aim to answer RQ1: ”What are people’s emotions and social roles like during COVID-19 home confinement?
+In this section, we aim to answer RQ1 ”What are people’s emotions and social roles like during COVID-19 home confinement?
 To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”.
-To achieve this, we analyze 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1.
-Furthermore, we explore the influence of COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
+To achieve this, we analyse 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1.
+Furthermore, we explore the influence of the COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
 4.1. The Perceived Valence, Arousal and Social roles
@@
 All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement.
-Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage 4 restrictions (strict home confinement) were lifted and steps for reopening commenced.
+Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage IV restrictions (strict home confinement) were lifted and steps for reopening commenced.
 Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges.
@@
 Next, we investigate the distribution of social roles and their correlation with emotion during home confinement.
-A social role refers to a set of norms, expectations, rules, beliefs and behaviors that are associated with a particular social situation (70).
-These roles can significantly influence users’ behaviors and the type of information they prefer to receive (3).
+A social role refers to a set of norms, expectations, rules, beliefs and behaviours that are associated with a particular social situation (75).
+These roles can significantly influence users’ behaviours and the type of information they prefer to receive (3).
 As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
@@
 work), with p-value less than 0.05.
-The above findings provides insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (31).
+The above findings provide insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (34).
 4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
@@
 I’m bored and sick of my housemates and it feels like time has stopped ’.
-Nevertheless, two felt the lockdown has a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
+Nevertheless, two felt the lockdown had a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
 (a) Productivity
@@
 Overall, while a minority of responses showed that lockdown has a positive impact on their lives (e.g., more time with family or gardening), most responses showed negative influences of lockdown on their valence, arousal and work-life balance.
-Besides the qualitative feedback from participants, we adopted Spearman correlation (23) to study the relationship between the reported Influence on Valence (from end-of-day survey) and the average perceived valence for that day (from ESM responses).
-In this way, we can determine whether the lockdown had a positive or negative effect on valence.
-Likewise, we went through the same process in terms of users’ arousal and social roles.
-For social roles, we correlate the reported Influence on wlb (‘wlf’ indicates work-life balance) to the percentage of the day spent on working (reported in ‘work’ or ‘both’ states).
-For example, if someone claimed that the lockdown significantly influenced their work-life balance (i.e., a score of 5, which becomes 3 after Likert transformation in the pre-processing), and they spent 50% of the day at work, then (50%, 3) would be a point in the correlation.
-5. Understanding App Usage Behaviors During COVID-19 Home confinement
+5. Understanding App Usage Behaviours During COVID-19 Home confinement
 In this section, we address RQ2 ”Is it possible to unsupervisedly extract user activity from app usage records?
@@
 Clustering.
-Next, AIGTS was conducted on all time series, resulting in a segmented list for each user.
+Next, AIGTS was conducted on all-time series, resulting in a segmented list for each user.
 To extract features from these segments, k-means clustering (51) was applied.
@@
 In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
-Figure 7. Average activities for all clusters
 Distribution of User Activities.
@@
 Conversely, the least frequent activities were Business Communication and Music, accounting for 1.22% and 1.65% of activities, respectively.
+Figure 7. Average activities for all clusters
 5.3. App Usage Correlation
@@
 Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day.
-The top correlation results for how different activities relate to usesr are displayed in Table 3.
+The top correlation results for how different activities relate to users are displayed in Table 3.
 We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
@@
 The correlation reaches a moderate level for older adults while weakening for the young age group.
-This might cause by the difference in working loads between young and older age groups.
+This might be caused by the difference in working loads between young and older age groups.
 Lifestyle negatively correlates with interruptions, especially for female and older adults.
@@
 Interruptions Lifestyle (-0.24) Communication Tools (0.42)
-Table 4. The computed features from smartphone usage behaviors
+Table 4. The computed features from smartphone usage behaviours
 Category Feature Description
@@
 Table 4 provides an overview of the features utilized in our research.
-Similar to previous mobile sensing studies (46; 76; 91), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
+Similar to previous mobile sensing studies (50; 81; 96), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
 To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences.
 Subsequently, we extract the features from these 90-minute segments.
-In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transforme them into sine and cosine values.
-Additionally, it is worth noting that app-usage patterns exhibit variations between weekdays and weekends (40; 46), prompting us to conduct separate analyses for these two categories.
+In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values.
+Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (44; 50), prompting us to conduct separate analyses for these two categories.
 Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night.
 Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile.
-For instance, the number of interacted notification is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’.
+For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’.
 Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
-To capture user activities related to app-usages, we employed our proposed method, AIGTS, as detailed in Section 5.2.
+To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 5.2.
 AIGTS allows us to extract both activities and their corresponding durations.
 Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval.
-In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combing two existing features.
-For example, the feature ‘num_communication’ is calculated by selecting and combing the notification sources.
+In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features.
+For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources.
 This composite feature counts the number of notifications originating the Business, Communication, or Social app categories.
@@
 6.2. Prediction Pipeline
-while emotion prediction can be addresed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (65; 93).
-We believe that utilizing regression is a better approach becuase valence and arousal labels are inherently ordinal data, where the relative order holds significance.
+while emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (70; 98).
+We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance.
 In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
@@
 Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’.
-Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicates ‘calm’ and 5 means ‘excited’.
+Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 means ‘excited’.
 To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
@@
 KNN is a widely utilized model known for its versatility across various domains.
-It works by finding the K nearest neighbors to a given data point and using their values to predict the target variable.
+It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable.
 RF and GB, on the other hand, are ensemble models based on decision trees.
@@
 This helps optimize the model’s performance.
-In the outer loop, similar to previous human-centered studies (30; 24), we divide the data into multiple groups, each group representing one participant.
+In the outer loop, similar to previous human-centred studies (33; 25), we divide the data into multiple groups, each group representing one participant.
 To ensure unbiased evaluation, we apply k-fold cross validation (93) where k=5 on all groups.
@@
 (c) RMSE for valence
-(d) MAE for arousal
-Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents a 95% confidence interval
+(d) RMSE for arousal
+Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents standard deviation.
 Baselines.
-Similar to previous human-centered studies (30; 87), we compare the proposed model with two baselines.
+Similar to previous human-centered studies (33; 92), we compare the proposed model with two baselines.
 The first baseline is the Mean baseline, which calculates the average score of user responses.
@@
 These factors may not be fully captured by mobile usage behaviours alone.
-Table 5. Prediction performance of the general models with different feature sets
-Prediction Result (RMSE/MAE)
+Table 5.
+Prediction performance of the general models for valence and arousal with different feature sets.
+The regression models are K-Nearest Neighbor (KNN), Random Forest (RF) and Gradient Boosting (GB), compared with Mean and Median baseline.
+The evaluation metrics are RMSE and MAE (with standard deviation).
+Prediction Result (RMSE(SD)/MAE(SD))
 Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
 All features
-Valence (29) 0.986/0.824 1.091/0.767 0.963/0.776 0.903/0.717 0.925/0.743
-Arousal (29) 1.113/0.884 1.146/0.784 1.088/0.844 1.005/0.766 1.084/0.756
+[…]
+[…]
 Features with $F$-statistic ¿= 5
-Valence (17) 0.986/0.824 1.091/0.767 0.956/0.769 0.904/0.717 0.987/0.725
-Arousal (14) 1.113/0.884 1.146/0.784 1.064/0.818 1.015/0.778 1.087/0.747
+[…]
+[…]
 Features with $F$-statistic ¿= 10
-Valence (13) 0.986/0.824 1.091/0.767 0.936/0.730 0.912/0.725 0.987/0.724
-Arousal (12) 1.113/0.884 1.146/0.784 1.064/0.819 1.011/0.774 1.076/0.745
+[…]
+[…]
 Table 5 displays the prediction results of the general models for valence and arousal.
@@
 Similar results were observed for the models utilizing the top 17 features.
-Regarding the arousal dimension, our results indicate that the RF regressor utilizing the all features achieved the best performance in terms of RMSE, with a value of 1.005.
+Regarding the arousal dimension, our results indicate that the RF regressor utilizing all features achieved the best performance in terms of RMSE, with a value of 1.005.
 This outperformed the Median baseline by 12.30%.
@@
 Additionally, it is worth noting that while the proposed valence models exhibited better prediction performance than baseline models for certain participants (e.g., ‘M02MMP’, ‘L10FGB’), there were participants who did not achieve satisfactory results.
-The possible reasons may be two-fold: firstly, individual differences play a crucial role in determining valence, and it may not be straightforward to accurately predict valence solely based on mobile usage behaviors for some participants.
+The possible reasons may be two-fold: firstly, individual differences play a crucial role in determining valence, and it may not be straightforward to accurately predict valence solely based on mobile usage behaviours for some participants.
 This highlights the complexity and subjectivity of valence and the need for a more nuanced understanding of individual emotional experiences; secondly, insufficient data availability may have limited the performance of the predictive models for some participants.
@@
 Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels.
-These findings differ slightly from the observations reported in (40), which found that app-usage patterns varied greatly vary by time and day.
+These findings differ slightly from the observations reported in (44), which found that app usage patterns varied greatly by time and day.
 The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends.
 This aligns with the findings presented in (18), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
-Above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement.
+The above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement.
 They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods.
@@
 7. Implications
-Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance.Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles.
-The findings from our study are significant in raising awareness about the importance of mental health not only during periods of home confinement but also in similar situations such as post-operative recovery, parental leave, aging, or limited mobility related to disability.
-These findings can also play a crucial role in the early detection and treatment of mental illnesses that may arise during these circumstances.
-Additionally, policymakers can utilize our research to develop appropriate strategies that minimize the negative effects of pandemic restrictive policies on mental health.
-To achieve this, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al.
+Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance.
+Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles.
+First of all, our findings highlight the importance of mental health during periods of home confinement.
+Home confinement is not limited to temporary pandemic adaptation but also the growing trend of working from home post-pandemic (40; 69) and other similar situations, such as post-operative recovery, parental leave, or limited mobility related to disability, or even mobile phone addiction.
+Secondly, our quantitative approach and regression models can contribute to tracking, early detection and treatment of mental illnesses during these circumstances.
+Our results also help policymakers to develop appropriate strategies that minimize the long-term negative effects of the pandemic on mental health and encourage people to build healthy digital habits (69).
+For example, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al.
 (12).
-In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (49).
-Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviors to those during the pandemic (53), underscoring the importance of studying people’s behaviors in high technology-reliance contexts.
+In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (53; 40).
+Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviours to those during the pandemic (57), underscoring the importance of studying people’s behaviours in high technology-reliance contexts.
 Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions.
@@
 For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
-In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviors of individuals under home confinement conditions, which has important implications for various fields.
+In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviours of individuals under home confinement conditions, which has important implications for various fields.
 For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation.
-Understanding how individuals cope with home confinement and the role that technology plays in their lives can help healthcare professionals or researchers develop effective interventions and design effective remote support programs to improve the overall quality of care for patients in similar situations.
-Furthermore, the findings from this study may have broader implications for our understanding of human behavior and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
+Rethinking about the future of work and life scenarios, where heavily rely on digital technologies.
+It is important to understand how people deal with being confined to their homes and how technology impacts their lives.
+This understanding can help researchers design smart environments, or assist healthcare professionals in developing effective remote support programs that enhance the quality of care for those in similar situations.
+Furthermore, the findings from this study may have broader implications for our understanding of human behaviour and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
 8. Limitations
@@
 These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4.
-In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage 4 restrictions was applied.
-While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (10; 14; 11).
+In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage IV restrictions was applied.
+While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (11; 15; 12).
 Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
@@
 Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses.
-It is worth noting that the delivery of end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience.
+It is worth noting that the delivery of the end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience.
 Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data.
@@
 9. Conclusion
-In this study, we delved into how human emotions, social roles, and mobile usage behaviors were affected during the COVID-19 home confinement period.
+In this study, we delved into how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period.
 Our proposed models yielded promising results in predicting human emotions during home confinement.
@@
 Notably, our findings revealed that time-related features emerged as the primary predictors for both valence and arousal.
-Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviors during home confinement is crucial.
+Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviours during home confinement is crucial.
 This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health.
-Moving forward, our research will expand to investigate users’ desktop usage behavior.
-By comparing the differences in usage behavior across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
+Moving forward, our research will expand to investigate users’ desktop usage behaviour.
+By comparing the differences in usage behaviour across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
 10. Acknowledgments
-We sincerely thank the anonymous reviewers.
+We sincerely thank the anonymous reviewers. We state that this manuscript has no relation to our prior publications (or concurrently submitted papers).

### v3

@@
 Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life.
-While essential in various scenarios, confinement within the home environment can profoundly influence psychological well-being and digital device usage.
-In this study, we delve into these effects, utilising the COVID-19 lockdown as a special case study to draw insights extending to various homebound situations.
-We conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks and analysed their emotions, well-being, social roles, and mobile usage behaviours.
-We extracted user activity from app usage records in an unsupervised manner, and experimental results revealed that app usage behaviours are effective indicators of emotional well-being in confined environments.
-Our research has great potential for developing supportive strategies and remote programs, not only for people facing similar medical isolation situations, but also for individuals in long-term home confinement, such as those with chronic illnesses, recovering from surgery, or adapting to permanent remote work arrangements.
+While essential in various scenarios, confinement within the home environment can profoundly influence mental well-being and digital device usage.
+Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of prolonged home confinement on mobile device usage.
+We conducted an in-situ study with 32 participants, analyzing three weeks of mobile usage data to assess emotional well-being and social dynamics in restricted environments.
+Our findings reveal that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation.
+This study highlights the potential for developing targeted interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
 Emotion, Home Confinement, Mobile Sensing, COVID-19 Lockdown, Social Roles, Well-Being
-copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXjournal: TOCHIccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
+copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
 1. Introduction
-Home confinement, defined as the restriction of individuals’ movements and activities within their homes (63), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (47; 17).
-Previous research indicated that home confinement may adversely affect mental well-being, leading to stress, anxiety, and depression (7; 78).
+Home confinement, defined as the restriction of movements within the home (52) due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives (38; 11).
+Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression (6; 66).
 For instance, Bartoszek et al.
-(7) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue.
-Understanding the impact of home confinement on mental well-being is essential for developing effective interventions to support individuals during prolonged periods of confinement.
-Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (78; 30).
-However, these approaches can be time-consuming and labour-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (32; 71)).
-Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being unobtrusively (73; 21).
-For instance, Rashid et al.
-(73) modelled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al.
-(21) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices.
-In home confinement scenarios, traditional mobile sensing methods, heavily relying on sensors like accelerometers and GPS to track physical behaviours, encounter unique challenges.
-These methods, which infer mental well-being from patterns such as visited locations, steps taken, and stationary periods (73), become less effective when individuals are confined at home.
-During home confinement, individuals’ interaction with their smartphones changes significantly.
-Rather than carrying their devices with them, people tend to leave their phones in one place.
-This shift in usage patterns diminishes the effectiveness of conventional metrics such as step counts and travel patterns, which are less indicative of mental well-being in a stationary context.
-Consequently, there is a shift towards analyzing more nuanced data types, such as application usage patterns, to better understand well-being under home confinement.
-Therefore, in this research, we focus on these finer-grained user activities to uncover deeper insights into the subtle changes in mental well-being during home confinement.
-Another challenge in studying the in-situ mental well-being of individuals confined to their homes is establishing a natural yet controlled environment.
-Home confinement situations vary significantly among individuals, influenced by different reasons and individual differences, making effective modelling challenging.
-Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement.
-The strict measures in place have created similar confinement conditions for all participants due to policy mandates.
-Additionally, the COVID-19 home confinement has led to an increasing reliance on digital technologies (68), with strategies such as stay-at-home and work-from-home being recommended worldwide (74).
-Recent research indicates that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (82).
-This reliance on technology provides an excellent way to understand people’s mental well-being during home confinement.
-In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during home confinement.
-Specifically, we focus on emotions and social roles, as these aspects significantly impact overall well-being (23; 65).
-To assess emotional states, we utilize the widely accepted Valence and Arousal dimensions within the Circumplex Model of Affect (79) to assess emotional states.
-Social roles, as defined by Role Theory (9), involve norms, expectations, rules and behaviours that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance.
-This, in turn, can have a significant impact on an individual’s overall well-being and mental health.
-Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behaviours.
-As a result, our research questions are:
-  • RQ1.
-What are people’s emotions and social roles like during COVID-19 home confinement?
-To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?
-  • RQ2. Is it possible to unsupervisedly extract user activity from app usage records? If so, how are these activities related to emotions, productivity and interruptions?
-  • RQ3. Can we model people’s emotion from mobile usage during home confinement? If so, what are the most significant types of features for prediction?
-To answer the research questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia.
-To answer RQ1, we analyzed patterns of perceived valence, arousal and social roles during COVID-19 home confinement using Experience Sampling Method (ESM) (8).
-We also investigated the influence of the COVID-19 lockdown on emotions, social roles, and work-life balance by examining the end-of-day survey data.
-To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day.
-Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviours for RQ3.
+(6) found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue.
+Understanding these impacts is essential for developing effective targeted interventions to support individuals during such periods.
+Traditionally, mental well-being during home confinement has been assessed using survey-based instruments (66; 22), which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias (23; 59).
+Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative (61; 15).
+While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation (61; 15), they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited (61).
+This study shifts focus to more granular data, such as app usage patterns, to provide deeper insights into emotional well-being during confinement.
+The diversity of individual circumstances complicates the study of home confinement.
+Factors such as the reason for confinement, personal health conditions, and social environment can vary greatly, complicating the creation of a standardized model.
+However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies (57; 70).
+This shared experience allows for a more uniform analysis of how mobile device usage reflects emotional states, social roles, and work-life dynamics during such periods.
+Espeically, while this study uses the lockdown as a descriptive case study (87), the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
+In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles, during COVID-19 lockdown, as a case study for home confinement.
+We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health (17; 54).
+By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect (67) and Social Role from the Role Theory (8), we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics.
+Our research seeks to address the following questions: RQ1.
+How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?
+RQ2.
+Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?
+To answer these questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia, involving 32 participants.
+We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) (7) and end-of-day surveys.
+We also proposed a novel method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records by identifying segments and transition times between apps during the day.
+The resulting insights offer a deeper understanding of how app usage during confinement reflects emotional well-being and social dynamics, with implications for developing digital interventions to support individuals in long-term home confinement.
 All scripts for this study are open source.
-111link to be shared after the double-blind review.
-In summary, our contributions are as below:
-  • We collected a comprehensive dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia.
-The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc.
-Their smartphone and desktop usage were recorded through an app running in the background.
-The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
-  • We analyzed participants’ perceived valence, arousal, and social roles during home confinement, and explored how lockdown affected well-being and work-life balance.
-We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
-  • We proposed an unsupervised activity recognition method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records.
-This approach autonomously identified 16 distinct user activities in an unsupervised manner, and we then explored the correlations between these user activities and human well-being.
-  • We conducted extensive experiments to predict human well-being in confined environments from smartphone usage behaviours, demonstrating the superior performance of our model.
-Additionally, we identified key factors contributing to well-being and discussed the primary predictors of human emotion during home confinement.
+[^1] Link will be shared after the double-blind review.
+In summary, our contributions are as follows:
+  • We collected a rich dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia.
+This dataset includes 502,485 records of users smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys.
+It provides a unique, real-world snapshot of emotional well-being and social dynamics during lockdown.
+  • We analyzed perceived valence, arousal, and social roles, revealing that the lockdown significantly impacted well-being, reducing emotional valence and increasing work hours due to interruptions or blurred work-life boundaries.
+These findings shed light on the complex emotional landscape during prolonged confinement.
+  • We introduced the App-based Information Gain Temporal Segmentation (AIGTS) method, an unsupervised technique for extracting user activities (e.g., video-watching, socializing, deep work) from app usage data.
+This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviours relate to emotional well-being.
+  • We demonstrated that mobile usage patterns can predict emotional well-being during home confinement.
+Our extensive experiments revealed key features that are most predictive of emotional states, offering new avenues for developing digital interventions to support individuals during isolation.
 2. Related work
-2.1. Impact of the Pandemic Restrictions on Human Well-being
-Pandemics and epidemics have played an essential role throughout human history (87).
-In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world.
-Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (87; 20; 12), including post-traumatic stress symptoms (76), depression (35), stress (26), emotional disturbance (97), irritability(48), low mood (48), confusion (16), anger (56) and anxiety-induced insomnia (26).
-Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (42).
-According to the World Health Organization  (94), the COVID-19 pandemic has severely impacted human well-being.
-The potential for physically sick or discomfort from COVID-19 is considered less of a concern than the psychological and social impact (38).
-A growing body of literature has explored the impact of COVID-19 on emotional and mental health (95; 22; 49; 46).
-Daly et al.
-(22) explored the effects of COVID-19 on mental health in the UK and found a significant increase in observed mental health problems.
-Fountoulakis et al.
-(28) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased by 10.40%.
-Khubchandani et al.
-(45) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after the initial COVID-19 lockdown in the USA.
-The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019.
-Killgore et al.
-(46) found that loneliness has increased during the COVID-19 pandemic, especially for those under lockdown restrictions.
-Despite the existing research on the impact of pandemic restrictions on human well-being, much of it is limited to specific regions and conducted during different phases of varying degrees of restrictive measures.
-This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being.
-In light of this, RQ1 focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage IV) were implemented.
-2.2. Impact of the Pandemic Restrictions on Smartphone Usage
-The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (89; 27; 4).
-Twitter even made an announcement to make working remotely a permanent option (13).
-As people perform their duties by staying at home or working from home, they increasingly turn to digital devices to communicate, interact and conduct job responsibilities (74).
-Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (82; 68).
-A recent global study (82) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (74).
-Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (74; 14; 5).
-Ratan et al.
-(74) pointed out that problematic smartphone usage affects the ‘reward system’ of the brain and subsequent harms will persist in the form of various emotional disorders even after the lockdown effects are lifted.
-Arora et al.
-(5) explored the problematic phone use and its impact on mental health during the COVID-19 lockdown, and found that ‘sadness’ and ‘anger’ were the dominant emotions in addictive Twitter users, while ‘valence’ was the dominating emotion among non-addictive users.
-Busch et al.
-(14) revealed that excessive mobile phone use may severely impact people’s mental health and human well-being.
-Especially during the COVID-19 lockdown, people’s mobile phone usage differs than normal, which increases the risk of overusing smartphones and affects their mental health (6).
-2.3. Investigating User Activities from App Sequences
-By analysing app sequences, researchers have gained insights into various aspects of individuals’ daily behaviours, including social roles  (2) and interruptiblity (66; 67).
-Deploying the approach ‘Application as a Sensor’, Anderson et al.
-(2) found that app sequences on smartphones are correlated with individuals’ private and work social roles.
-Similarly, Okoshi et al.
-(66; 67) utilized app sequence data and machine learning techniques to predict breakpoints in activity for intelligent notification delivery.
-These studies focus on identifying breakpoints between apps to understand interruptibility.
-However, no previous research has used breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
-Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period.
-However, those approaches fail to capture the nuances and transitions between different app activities, limiting understanding of user behaviours.
-Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data.
-Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information about user behaviours and mental well-being.
-With this in mind, RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
-2.4. Inferring Human Well-being Using Mobile Sensing Technologies
-The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (55) and psychological stages, e.g., valence (41), mood (60; 98; 59), engagement (33; 39; 31), social anxiety (21; 73), depression (91; 88), etc.
-Most studies integrate data from various sources for building comprehensive prediction models.
-For instance, Gao et al.
-(33) predicted emotional, cognitive and behavioural student learning engagement in an Australian high school using wearable sensors and environmental sensors.
-Wang et al.
-(91) measured depression of undergraduate students at Dartmouth College using passive mobile sensing data and wearable data.
+2.1. Impact of the Home Confinement on Mental Well-Being
+Home confinement, whether due to edical conditions, disability, or societal events like pandemics, can significantly affect mental well-being.
+Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness (76; 14; 10).
+The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger (64; 39).
+These effects may persist even after restrictions are lifted (31).
+The COVID-19 pandemic underscored these effects on a global scale.
+The World Health Organization (84) reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks.
+Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread (85; 16; 37).
+For example, Khubchandani et al.
+(36) found depression (39%) and anxiety (42%) surged in the U.S.
+post-lockdown.
+Similarly, Fountoulakis et al.
+(20) reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece.
+Loneliness was also a major issue, especially among those in long-term isolation (37).
+Despite extensive research on the mental well-being impacts of various forms of home confinement, much of it is geographically limited or reflects different phases of isolation.
+This variability makes it challenging to fully understand the broader, long-term psychological effects of home confinement in diverse contexts.
+To address this gap, RQ1 examines how emotions and social roles are experienced during COVID-19 home confinement, using self-reports and app usage patterns to infer these dynamics.
+2.2. Digital Device Usage During Home Confinement
+During periods of home confinement, digital devices play a crucial role in maintaining communication, accessing information, and fulfilling work and social roles.
+With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities (32).
+For instance, Ratan et al.
+(62) found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
+While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behaviour under prolonged confinement.
+With workplaces and schools shifting to remote environments, the demand for digital communication tools surged (79; 19; 5).
+Video conferencing platforms, social media, and messaging apps saw massive increases in usage (70; 57), and overall digital device use skyrocketed.
+A global survey found 70% of internet users reported increased smartphone usage during the lockdown (70), emphasizing how integral smartphones became in managing both work and personal life.
+Similarly, Majumdar et al.
+(45) observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown.
+Jonnatan et al.
+(33) also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
+In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication (43; 29).
+Studies suggest that behaviours such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase (46).
+This highlights the importance of understanding digital device usage patterns in today’s technology-reliant environments, particularly in contexts of home confinement.
+2.3. Inferring Mental Well-Being Using Mobile Sensing Technologies
+Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood (49; 88; 48), engagement (24; 28; 2), and mental health issues like depression and anxiety (15; 61; 81).
+Many studies have integrated data from multiple sensors to improve prediction accuracy.
+For example, Gao et al.
+(24) used wearable and environmental sensors to predict student engagement, while Wang et al.
+(81) tracked depression using passive mobile sensing and wearables.
 Jaques et al.
-(41) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behaviour data collected from mobile devices.
-Morshed et al.
-(60) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
-While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behaviour, e.g., physical activity, communication patterns, and location.
-For example, Hintze et al.
-(37) analyzed mobile device usage characteristics such as session length, interaction frequency and daily usage in locked/unlocked state with respect to location context and diurnal pattern, to gain insights into how people use their smartphones and tablets.
-Madan (55) et al.
-designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
-Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (73; 88; 21), mood (98; 59), emotion (86), social roles (2) and interruptibility (66; 67).
-Zhang et al.
-(98) proposed an Android app to detect compound emotion utilizing smartphone usage data and Meegahapola et al.
-(59) explored the effect of geographical diversity on mood inference.
-Tlachac et al.
-(88) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine-learning models to predict anxiety and depression.
-Tag et al.
-(86) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings.
-While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement.
-Unlike daily life situations, individuals’ behaviours may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working.
-Although a few studies (88; 64) have examined psychological states using mobile sensing during the COVID-19 pandemic, these works largely focus on broader environments rather than strictly home confinement.
-Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, ageing or disability-related limited mobility.
-Therefore, we aim to address RQ3 by exploring the possibility of modelling emotions during home confinement and identifying significant features for prediction.
-3. Data Collection
-We conducted a field study in Melbourne, Australia, for three weeks, from 24 August 2020 to 13 September 2020.
-The study has been approved by the Human Research Ethics Committee at the authors’ university.
-Prior to the data collection, we obtained informed consent from all participants, ensuring that they were aware of the study’s purpose and that the dataset would be publicly released.
-The details about the background, participants, and collected data will be described in this section.
-3.1. Background
+(30) modelled student mood using physiological signals and mobile behaviour.
+While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviours like physical activity, communication patterns, and location (25).
+Studies have shown that mobile usage data can predict psychological states such as anxiety (61; 77), mood (88; 48), and emotion (75).
+For instance, Zhang et al.
+(88) developed an app to detect compound emotions from smartphone usage, and Tlachac et al.
+(77) used call and text logs to predict anxiety and depression during the pandemic.
+Research on app sequences has also provided insights into daily behaviours, such as social roles (3) and interruptibility (55; 56).
+Anderson et al.
+(3) demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al.
+(55; 56) used app sequences to predict activity breakpoints for intelligent notification delivery.
+However, most studies have focused on identifying breakpoints rather than modeling user activity or predicting contextual variables.
+Additionally, earlier studies often relied on basic aggregation methods, which fail to capture transitions between activities.
+Thus, RQ1 also explores emotions and social roles by analyzing app usage patterns.
+While mobile sensing technologies have shown potential in predicting well-being, gaps remain in understanding how these tools can be applied specifically during home confinement.
+Home confinement introduces unique behaviours—such as increased reliance on digital devices and altered perceptions of time—that differ from daily life.
+Although some studies (77; 53; 74) have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement.
+Understanding this unique setting could provide valuable insights for other confinement scenarios, such as post-operative recovery, parental leave, or aging.
+Therefore, RQ2 investigates how mobile sensing can model emotions during home confinement and identify key features for prediction.
+3. Study Overview
+We conducted a three-week field study in Melbourne, Australia, from 24 August 2020 to 13 September 2020, during the strictest Stage IV COVID-19 restrictions.
+This study use Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work.
+The following section details the background, participants, and data collected.
+3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia
 Figure 1. The periods of six lockdowns and our data collection
-To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (58; 85).
+To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (47; 73).
 Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1).
@@
 On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions.
-During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (74; 82).
-This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviours of individuals during home confinement.
-The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
-(a) Age and gender
-(b) Smartphone and desktop
-Figure 2. Demographic and device usage information of participants
-3.2. Participants
+During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (62; 70).
+3.2. Procedures
 Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers.
-In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, std = 8.95).
-The age distribution of participants across genders can be found in Figure 2.
-It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample.
-While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behaviour.
-Rather, we focus on individuals’ psychological states and smartphone usage behaviours during COVID-19 home confinement.
-Therefore, we believe that the potential oversample threat to the validity of our results is minimized.
-However, it is still important to note that our findings are specific to this particular population subset and may not be necessarily generalizable to other populations.
-Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
+In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, SD = 8.95).
+Of these, 27 participants provided smartphone data, while 13 contributed desktop data.
+Recruitment was primarily through social media platforms Facebook and Discord, potentially leading to an over-representation of users from these platforms.
+However, as our study focused on psychological states and digital device usage during COVID-19 home confinement, not social media behaviour, we believe this does not significantly impact the validity of our findings.
+Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps(27; 80) on their smartphones and desktops.
+These apps, designed to record usage behaviors and enable experience sampling, were optimized for minimal computational, energy, and storage consumption.
+The apps used Experience Sampling Method (ESM) to prompt participants with a short questionnaire every 90 minutes during active device use between 7 am and 10 pm.
+To minimize participant burden, surveys were only triggered during active use.
+Additionally, the event-based ESM was triggered after 10 minutes of interaction, with follow-ups every 30 minutes.
+At day’s end, participants received an end-of-day survey via email at 7:30 pm.
 After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time.
-Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage.
-Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants.
-Analysis of desktop usage behaviours will be explored further in future work.
-3.3. Procedures
-Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (36) on their smartphones and desktops, respectively.
-These applications were specifically designed to record usage behaviours and enable background sensing and experience sampling, drawing inspiration from Van et al.
-(90).
-Both applications were optimized to minimize their consumption of computational power, energy, and storage resources.
-The applications utilized ESM, popping up a short questionnaire to participants every 90 minutes when their devices were actively in use between 7 am and 10 pm.
-This helps to obtain unbiased samples of participants’ activities, which served as the ground truth for our predictive models.
-Furthermore, an event-based ESM was triggered on the user’s smartphone after a minimum interaction of 10 minutes, with subsequent event-based ESMs occurring every 30 minutes.
-At the end of each day, participants received an end-of-day survey via email at 7:30 pm.
-This survey aimed to gather reflections on their well-being and the impact of the lockdown on their daily experiences.
-This provides valuable context and additional data beyond the ESM questionnaires.
-Overall, we collected 502,485 recordings for smartphone usage, 1,749 ESM responses, and 265 end-of-day responses.
-Table 1. An overview of the collected data
+Privacy, ethics, and safety.
+Our field study was designed and conducted after carefully reviewing potential risks and benefits, in compliance with the city’s COVID-19 safety regulations.
+The study has been approved by the Human Research Ethics Committee at the authors’ university.
+To ensure privacy, all personal identifying information was anonymized before data analysis.
+Participants’ consent was obtained prior to data collection, with a clear explanation of the study’s purpose and the nature of the data being collected.
+We also implemented strong data security measures, including encryption of sensitive data, to further protect participants’ privacy.
+Participants were informed of their right to withdraw at any time without penalty, and their data would be excluded from the study upon request.
+Table 1. Overview of the collected data records across devices
+Category Device Total Records Participants Data Used for RQs
+App Both 502,485 32 RQ1
+Desktop 261,911 13 /
+Smartphone 240,574 27 RQ2
+ESM Both 1,749 28 RQ1
+Desktop 227 10 /
+Smartphone 1,522 25 RQ2
+Ene-of-Day Survey Both 265 25 RQ1
+Table 2. Detailed overview of data categories, items, and response options
 Category Items Questions / Descriptions Options
-Smartphone Usage Application usage Foreground application name, package identifier, category N/A
+Device Usage Application usage Foreground application name, package identifier, category N/A
 Notifications Notification arrival time, content (hashed), sender (hashed), relationship N/A
@@
 Difference How was today different from yesterday? Text (optional)
-3.3.1. Smartphone usage data
-During the data collection, the background services of smartphones and desktops continuously tracked the following information: application usage (using Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (using Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (i.e.
+3.3. Collected Data
+The data collection period spanned 21 days.
+From the 32 recruited participants, we collected 502,485 recordings of smartphone/desktop usage, 1,749 ESM responses, and 265 end-of-day surveys.
+Table 1 provides an overview of the collected data across devices.
+To address RQ1, we used both desktop and smartphone data to explore how well-being was during COVID-19 home confinement.
+For RQ2, due to the limited number of participants contributing desktop data, we focused exclusively on smartphone data and its corresponding ESM responses.
+Smartphone/desktop usage data.
+During the data collection, background services of smartphones and desktops continuously tracked the application usage (via Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (via Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (only for smartphone, i.e.
 screen/battery status, ringer modes, location updates and physical activities).
-Specifically, the mobile app usage behaviour was recorded every time the user opened the app.
-For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none).
-Table 1 displays an overview of the collected data for smartphone usage, ESM and end-of-day survey.
-In total, we have collected 502,485 recordings for smartphone usage data from 27 participants.
-3.3.2. ESM responses
-There are some popular questionnaires to measure well-being (43; 10), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (10) was employed in this research.
-In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users.
-Especially, the item ‘valence’ was rated with a 5-point Likert scale from ‘unhappy’ to ‘happy’, and the item ‘arousal’ was also rated with a 5-point Likert scale from ‘calm’ to ‘excited’.
-Both valence and arousal were used to represent the users’ emotions based on the circumplex model (79).
-Understanding the social role, interruptiblity and task help us understand how people manage their time and attention on the digital devices (3).
-The detailed items of the ESM can be found in Table 1 and the screenshot of the ESM on users’ smartphones was shown in Figure 3(a).
-Overall, we have collected 1,749 ESM responses from 29 participants.
-Of these, 1,522 responses were obtained from 25 participants using smartphones (2 participants did not contribute any ESM responses), while the remaining 227 responses were from 10 participants using desktops.
+Specifically, smartphone app usage was recorded each time an app was opened, while desktop usage was tracked by polling the active application every second.
+For notifications, data captured included arrival time, hashed content (for notification length), sender (private, work, both, or none), and the participant’s relationship to the sender (family, friend, work, or none).
+Table 2 displays an overview of the collected data for digital device usage, ESM and end-of-day survey.
+ESM responses.
+There are some popular questionnaires to measure well-being (34; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research.
+In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and tasks from users (see Table 2).
+Specifically, the measurement of items valence and arousal were adapted from SAM  [^4] Compared to the original SAM questionnaire, we added a time constraint, asking participants to describe their feelings over the past hour.
+and rated with a 5-point Likert scale, from unhappy to happy for valence and from calm to excited for arousal.
+These measures were used to represent the users’ emotions based on the circumplex model (67).
+The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework (4).
 (a) ESM
@@
 Figure 2. The screenshots for the ESM and End-of-day survey
-3.3.3. End-of-day survey
-As shown in Figure 3(b), the end-of-day surveys were optional and included ten items to understand participants’ well-being during the lockdown such as productivity, valence, energy, and work-life balance.
-In addition, we surveyed how participants feel today compared to yesterday.
-This information could describe the dynamics of people’s experience during the COVID-19 lockdown.
-In the end-of-day survey, each item was rated with a 5-Likert scale or answered with free text input (see Table  1).
-Specifically, item 1 and item 2 were rated from 1 to 5 which indicates ‘strongly agree’, ‘somewhat agree’, ‘neither agree nor disagree’, ‘somewhat disagree’, ‘strongly disagree’, and items 3, 4, 5, 9 were rated from 1 to 5 which indicates ‘a great deal’, ‘a lot’, ‘a moderate amount’, ‘a little’, ‘not at all’.
-In total, we collected 265 responses from 25 participants during the data collection period.
-4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
-In this section, we aim to answer RQ1 ”What are people’s emotions and social roles like during COVID-19 home confinement?
-To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”.
-To achieve this, we analyse 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1.
-Furthermore, we explore the influence of the COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
-4.1. The Perceived Valence, Arousal and Social roles
-The distribution of perceived valence, arousal, and social roles from the ESM responses is depicted in Figure 4.
-On average, participants reported a valence level of 3.46 (SD = 1.01) and an arousal level of 2.72 (SD = 1.11).
-Usually, the relatively high valance and low arousal refer to a state of calmness, relaxation and peace.
-This could be attributed to various factors during the COVID-19 home confinement.
-One possible reason is that people may feel relieved to be safe at home, and cherish the opportunity to spend more quality time with their family and loved ones, away from the potential risks of infection outside.
-Furthermore, the absence of the usual stressors of daily life, such as commuting, work pressures, and social obligations, may be contributing to a general sense of calm and well-being.
-Additionally, some people may have found comfort in the sense of community and solidarity that emerged during the pandemic, as people came together to support one another through difficult times.
-All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement.
-Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage IV restrictions (strict home confinement) were lifted and steps for reopening commenced.
-Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges.
-For instance, prolonged periods of confinement may lead to feelings of boredom, loneliness, and frustration.
+End-of-day survey.
+As shown in Figure 2(b), the end-of-day surveys were optional and included ten items designed to understand participants’ well-being during the lockdown, focusing on aspects such as productivity, valence, energy, and work-life balance.
+Additionally, we surveyed participants on how they felt today compared to yesterday.
+This information helps describe the dynamics of people’s experiences during the COVID-19 lockdown.
+The end-of-day survey questions were tailored to capture the unique context of our study and were not sourced from existing questionnaires.
+Each item in the end-of-day survey was rated using a 5-point Likert scale or answered with free text input (see Table  2).
+Specifically, items 1 and 2 were rated from 1 to 5, indicating strongly agree, somewhat agree, neither agree nor disagree, somewhat disagree, and strongly disagree.
+Items 3, 4, 5, and 9 were rated from 1 to 5, indicating a great deal, a lot, a moderate amount, a little, and not at all.
+4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement (RQ1)
+In this section, we address RQ1 ”How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?”.
+First, we analyse 1522 ESM responses collected via smartphones from 25 participants.
+Next, we explore the impact of the COVID-19 lockdown by examining the end-of-day survey.
+Additionally, we explore users’ app usage behaviours through unsupervised activity recognition, identifying common activity clusters and their relationship to the perceived psychological metrics.
+4.1. The Perceived Valence, Arousal and Social Roles During Covid-19 Home Confinement
+Figure 3 illustrates the distribution of perceived valence, arousal, and social roles.
+On average, participants reported a valence of 3.46 (SD = 1.01) and arousal of 2.72 (SD = 1.11), indicating a generally calm and relaxed emotional state.
+This likely reflects feelings of safety and relief from the external stressors associated with the pandemic, such as the risks of infection.
+The absence of daily stressors—like commuting, work pressure, and social obligations—may also contribute to this sense of well-being.
+Furthermore, the solidarity and support networks that emerged during the pandemic likely provided additional emotional comfort.
+However, it is important to note that our data collection occurred during a three-week period as the COVID-19 Stage IV restrictions were lifting.Had the confinement continued longer, participants likely would have reported more negative emotions, such as boredom or frustration, as they adjusted to the new normal.
+Next, we investigate the role of social roles in shaping emotional experiences during home confinement.
+A social role refers to the expectations, norms, and behaviours associated with particular social contexts (63).
+These roles usually influence users’ behaviours and the type of information they prefer to receive (4).
+As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
+We further explore how emotions vary across different social roles.
+Figure 3(c) shows that participants reported the highest valence and lowest arousal when engaged in private-related roles, compared to work-related or hybrid roles.
+This suggests that personal activities, which are often less stressful and more enjoyable, foster a greater sense of well-being.
+Conversely, work-related and hybrid roles were associated with higher levels of stress and pressure.
+Welch’s t-tests (89) confirm these findings, showing significant differences in valence between private-related and the other two roles (i.e., private vs.
+both, private vs.
+work), with p-value ¡ 0.05.
+The findings offer insights into how individuals can structure their daily routines to boost positive emotions during home confinement.
+Interventions aimed at improving well-being in such situations may benefit from considering the social roles individuals are engaging in (26).
 (a) Valence and arousal
@@
 Figure 3. The distribution of perceived valence, arousal, social roles and their relationship
-Next, we investigate the distribution of social roles and their correlation with emotion during home confinement.
-A social role refers to a set of norms, expectations, rules, beliefs and behaviours that are associated with a particular social situation (75).
-These roles can significantly influence users’ behaviours and the type of information they prefer to receive (3).
-As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
-Since social roles play a crucial role in shaping individuals’ well-being, we investigate the emotions of users in different social roles, as shown in Figure 4(c).
-We found that, when engaging in private-related roles, participants had the highest levels of valence and lowest levels of arousal compared to work-related roles and hybrid roles.
-This may be because private-related roles involve personal activities which can provide a sense of comfort and happiness, while work-related roles or hybrid roles may involve stress, deadlines, and pressure to perform.
-The results of Welch’s t-tests (99) further support these findings, indicating that there is a significant difference in valence levels between private-related roles and the other two roles (i.e., private vs.
-both, private vs.
-work), with p-value less than 0.05.
-The above findings provide insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (34).
-4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
-In total, 54.3%, 51.3% and 54.7% of end-of-day responses displayed that the lockdown has influenced their valence, energy, and work-life balance to some extent, as shown in Figure 5.
-For the influence of lockdown on valence, 13 participants reported that it had a negative effect, claiming they ‘felt quite anxious and sad’ as ‘staying at home has made me a little sad’.
-They complained that they were unable to see friends (9 responses), see families (5 responses), go outside (10 responses), and feel bored (6 responses), e.g., ‘I thought I might be able to see friends, but I did not’, ‘I want to go outside, to the park’, ‘I’m missing my family’, ‘It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long.
-I’m bored and sick of my housemates and it feels like time has stopped ’.
-Nevertheless, two felt the lockdown had a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
+4.2. The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
+End-of-day survey responses reveal that the lockdown had a noticeable impact on participants’ valence, arousal, and work-life balance.
+As shown in Figure 4, 54.3%, 51.3%, and 54.7% of responses indicated that the lockdown influenced valence, energy, and work-life balance, respectively.
+In terms of valence, 13 participants reported negative effects, describing feelings of anxiety, sadness, and boredom.
+Common complaints included missing social interactions (9 responses), the inability to go outside (10 responses), and a sense of boredom (6 responses).
+For instance, ”I thought I might be able to see friends, but I did not”, ”I want to go outside, to the park”, ”I’m missing my family”, ”It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long.
+I’m bored and sick of my housemates and it feels like time has stopped”.
+On the other hand, two participants reported positive effects, such as feeling happier or using the extra time for physical activity and seamless work transitions.
 (a) Productivity
@@
 Figure 4. The distribution of productivity, interruption, and influence during the lockdown
-Regarding the impact of lockdown on arousal, 11 participants claimed that the impact was negative, reporting feeling tired (11 responses), being lazy (5 responses) and losing motivation (3 responses), e.g., ‘I don’t feel as motivated to do productive things’, ‘Emotionally tired, perhaps as a result of the lockdown’.
-However, two participants claimed the lockdown gave them ‘more energy’.
-For the influence on work-life-balance, most participants reported that they faced more interruptions or distractions (14 responses), had difficulty separating work and life (11 responses), spent more time on work due to interruptions or blurring of personal and work boundaries (15 responses), e.g., ‘I worked a bit more than I should’ve’, ‘Everything blends together, separating work and life is really difficult’, ‘Continuous interruptions, even my cat interrupts me’, ‘Need to work longer hours due to interruptions’.
-Overall, while a minority of responses showed that lockdown has a positive impact on their lives (e.g., more time with family or gardening), most responses showed negative influences of lockdown on their valence, arousal and work-life balance.
-5. Understanding App Usage Behaviours During COVID-19 Home confinement
-In this section, we address RQ2 ”Is it possible to unsupervisedly extract user activity from app usage records?
-If so, how are these activities related to emotions, productivity and interruptions?”.
-First, we introduce the process of app categorization.
-Next, we conduct unsupervised activity recognition from app sequences and display the average activities for all clusters.
-In this end, we explore the correlation between user activities and perceived psychological metrics.
-Table 2. App categories and top three frequently used apps
-Upper Category Category Top App Second App Third App
+Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., ”I don’t feel motivated to do productive things”, ”Emotionally tired, perhaps as a result of the lockdown”.
+However, two participants noted that they felt more energized during the lockdown.
+In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as ”I worked a bit more than I should’ve”, ”Everything blends together, separating work and life is really difficult”.
+These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities.
+While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
+4.3. Understanding App Usage Behaviours During COVID-19 Home Confinement
+Table 3. App categories (referred to the Google Play Store categories) and top three frequently used apps.
+Upper Category Subcategory Top App Second App Third App
 Business Business Teams LinkedIn Zoom
@@
 Personalization One UI Home OnePlus Launcher TouchWiz home
-5.1. App Categorization
-We categorized the apps using historical usage logs.
-For Android apps, we referred to the categories listed in the Google Play Store, excluding uncategorized apps.
-Windows apps were manually sorted into the same categories as Google Play Store apps.
-Less frequently used apps, with a cumulative usage time of less than one hour across all participants, were excluded.
-Overall, we recorded 28 app categories and consolidated them into 8 upper-level categories: Business, Communication, Entertainment, Productivity, Reading, Social, Tools and Others.
+4.3.1. App Categorization
+Apps were categorized based on historical usage logs.
+For Android apps, we used the categories from the Google Play Store, excluding uncategorized apps.
+Desktop applications were manually assigned to equivalent categories from the Google Play Store.
+Apps with less than one hour of cumulative usage across all participants were excluded.
+In total, we identified 28 app categories, which were consolidated into 8 high-level categories.
 A summary of the top three most frequently used apps in each category is provided in Table 3.
-Notably, Chrome is categorized as a Communication app due to the scarcity of browser apps in the Google Play Store.
-Categorizing desktop apps within the same categories used for mobile apps proves challenging, as they are often used for work and software development (e.g., Pycharm, WinSCP etc).
-In light of this, we grouped several development and utility apps under the Tools Category.
-Overall, the top 10 Tools apps for both desktop and mobile include: Gboard, Explorer, Samsung Keyboard, Always on Display, WinSCP, Pixel Launcher, Acrobat, pychar64, Google and Samsung keyboard.
-It is worth mentioning that developer tools such as Pycharm and WinSCP, as well as the utility Explorer, fall under this category.
-5.2. Unsupervised Activity Recognition From App Sequences
+Notably, Chrome was categorized under Communication, as the Google Play Store has few dedicated browser apps.
+Categorizing desktop applications posed a challenge, as many are used for work or software development (e.g., PyCharm, WinSCP).
+These were grouped under the Tools category, along with utility apps like Explorer.
+4.3.2. Unsupervised Activity Recognition From App Sequences
+To identify distinct user activities from app usage, we propose an unsupervised approach that combines segmentation and clustering.
+First, we segment the app usage sequences to capture shifts in activity, then cluster these segments to identify common patterns of behaviour.
 Segmentation.
-To study how the application usage is related to the aforementioned variables, we propose App-based Information Gain Temporal Segmentation (AIGTS), which focuses on differentiating activities within app usage.
-Figure 6 provides an illustration of how activities are extracted from app usage data.
-AIGTS draws inspiration from Information Gain Temporal Segmentation (IGTS) (80), an unsupervised method for segmenting multivariate time series data.
-IGTS works by identifying which splits in the multivariate time series minimize the weighted entropy across the series.
-The variables within the multivariate time series data corresponded to whether an app of a particular category was used.
-Unlike IGTS focusing solely on the time series, AIGTS aims to identify different segments and transition times between apps throughout the day.
-It is capable of detecting when participants change their app usage patterns, such as transitioning from using Email and Browsers (both categorized under Communication) to browsing Reddit (News and Media category).
-This method allows for segmentation at any level of granularity desired, ranging from half a day to as little as half a minute for application activity.
-The granularity of segmentation is determined by the variable $k$, which represents the number of splits made in the time series.
-To ensure that segments are captured at a consistent level of granularity, we propose the following steps for calculating the optimal $k$ in AIGTS:
-  1. (1)
-    Calculate the knee point of the activity data for each day;
-  2. (2)
-    Identify the average number of segments per minute for all days based on the knee point;
-  3. (3)
-    Compute the median average segments per minute for all users;
-  4. (4)
-    Multiply this median by the number of minutes per day (which varies from the first log to the last log) and choose this value as $k$.
-These procedures guarantee that the average length of each segment will be consistent across all days.
-Figure 6. An illustration of unsupervised activity recognition using app usage data
+To investigate the relationship between app usage and various factors, we propose App-based Information Gain Temporal Segmentation (AIGTS), a method that focuses on identifying different activities within app usage.
+Figure 5 illustrates how activities are extracted from app usage data.
+AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) (68), an unsupervised method for segmenting multivariate time series.
+IGTS identifies splits that minimize the weighted entropy across the multivariate time series, where each variable corresponds to whether a specific app category was used.
+Unlike IGTS, which focuses solely on the time series, AIGTS detects activity transitions, such as switching from Email or Browser usage (both in the Communication category) to browsing Reddit (in the News and Media category).
+This method can segment app usage data at various granularities, ranging from half a day to just 30 seconds.
+The level of segmentation is determined by the variable $k$, which represents the number of splits made in the time series.
+To ensure consistency in segment granularity across days, we compute the optimal $k$ using the following steps: (1) Identify the knee point of the activity data for each day; (2) Calculate the average number of segments per minute for all days based on the knee point; (3) Compute the median segments per minute across all users; (4) Multiply this median by the number of minutes each user spent on their device per day and select this value as $k$.
+These steps ensure that the average segment length remains consistent across all days.
+Figure 5. An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
 Clustering.
-Next, AIGTS was conducted on all-time series, resulting in a segmented list for each user.
+Next, AIGTS is applied to the all-time series, resulting in a segmented list for each user.
 To extract features from these segments, K-means clustering (41) was applied.
@@
 In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
-Distribution of User Activities.
-Following the segmentation and clustering, we uncovered 16 clusters.
-The average duration of these user activities was 18 minutes.
-Figure 7 displays average activities for all clusters, where the average activity represents the average distribution of app categories appearing in each segment within the cluster.
-These clusters were manually named to describe the distribution of average activities.
-The $x$-axis displays the app categories that frequently appear in clusters, with infrequently appearing categories omitted.
-Notably, the cluster Other includes all uncategorizable segments and has roughly equal means across multiple categories.
-Additionally, the cluster Idling is used for Personalization apps because Personalization apps are typically launchers (see table 2), and indicate that the user is sitting on their device’s home screen.
-Moreover, the most frequent user activities during the COVID-19 lockdown were Just Communication and Idling, accounting for 17% and 14.60% of activities, respectively.
-Conversely, the least frequent activities were Business Communication and Music, accounting for 1.22% and 1.65% of activities, respectively.
-Figure 7. Average activities for all clusters
-5.3. App Usage Correlation
-Furthermore, we investigate the correlations between user activities with their valence, arousal, productivity, and interruptions.
-To understand whether any particular activity $A$ was correlated with valence, we ran a Spearman correlation (24) between valence scores and the amount of time spent in user activity $A$ in the previous hour.
-The same goes for the arousal, productivity, and interruptions, but the productivity and interruptions vary in duration from 1 hour to 24 hours, as these are calculated from the end-of-day survey and correspond to the value for a full day.
-Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day.
-The top correlation results for how different activities relate to users are displayed in Table 3.
-We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
-As shown in Table 3, the correlations are weak with most activities, but few activities have moderate correlations.
-Tools related activities were positively correlated with overall valence and for both age groups.
-Particularly, Idle Tools have a relatively stronger correlation with valence in older adults, indicating an increase in the mood at work (Idle Tools include phone home screens and phone keyboards).
-Surfing by women and young age group are negatively correlated to valence, indicating reading news or surfing (e.g., reading COVID-19-related news) are negatively related to their valence.
-Productivity was negatively correlated with Communication Tools (e.g., keyboards and chat clients), and positively correlated with Business Communication (i.e., both Communication and Business apps such as Microsoft Teams).
-It can be seen that while both are Communication applications, whether or not they are used for work may imply productivity levels.
-Interruptions can be seen as the inverse of productivity, as it has positive correlation with Communication Tools, but the value is higher.
-The correlation reaches a moderate level for older adults while weakening for the young age group.
-This might be caused by the difference in working loads between young and older age groups.
-Lifestyle negatively correlates with interruptions, especially for female and older adults.
-Table 3. Top correlations for all variables
+Figure 6. Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
+Overview of User Activities.
+Following segmentation and clustering, we identified 16 distinct clusters representing user activities, with an average activity duration of 18 minutes.
+Figure 6 presents a heatmap showing the proportion of app subcategories in each cluster.
+The $x$-axis denotes the app categories (categories that infrequently appear in each segment are omitted), while the $y$-axis represents the clusters (manually named such as Deep Work, Hybrid Work).
+The color intensity indicates the relative contribution of each app category to the cluster’s overall activity distribution.
+The Device Idle cluster, corresponding to Personalization apps (e.g., launchers, as shown in Table 3), represents periods when users remain on their device’s home screen.
+We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the COVID-19 home confinement, accounting for 17% and 14.6% of all activities, respectively.
+These activities are characterized by high proportions of communication-related apps and personalization apps, suggesting users spent considerable time on their devices in passive states or communicating with others.
+In addition, clusters such as Work-Related Communication and Hybrid Work exhibited higher proportions of productivity and business apps.
+These clusters likely reflect professional tasks and remote work activities, which were common during the lockdown period.
+The differentiation between Light Work and Deep Work clusters highlights variations in the intensity and type of work-related app usage.
+4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
+.
+We further investigate how user activities and app usage behaviours correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions.
+To assess these correlations, we performed Spearman Rank Correlation (18) tests between the time spent on different activities/apps and the scores for each dimension.
+Specifically, valence and arousal were examined for theirr relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
+Table 4. Top correlations between app usage/user acitivty and well-being dimensions.
 Groups Variable Lowest correlation $\rho$ Highest correlation $\rho$
-All Valence Work and Write (-0.07) Idle Tools (0.11)
-Arousal Business (-0.05) Social Tools (0.15)
-Productivity Communication Tools (-0.24) Business Communication (0.18)
-Interruptions Lifestyle (-0.15) Communication Tools (0.29)
-Male Valence Just Communication (-0.08) Business Communication (0.10)
-Arousal Idling (-0.12) Just Communication (0.11)
-Productivity Communication Tools (-0.15) Business Communication (0.34)
-Interruptions Video Watching (-0.23) Communication Tools (0.18)
-Female Valence Surfing (-0.17) Other (0.13)
-Arousal Work and Write (-0.17) Social Tools (0.21)
-Productivity Communication Tools (-0.30) Lifestyle (0.16)
-Interruptions Lifestyle (-0.33) Idling (0.41)
-Age 18-24 Valence Surfing (-0.13) Just Tools (0.09)
-Arousal Idle Tools (-0.17) Social Tools (0.19)
-Productivity Communication Tools (-0.21) Work and Write (0.20)
-Interruptions Social Tools (-0.10) Communication Tools (0.16)
-Age $\geq$25 Valence Just Communication (-0.12) Idle Tools (0.23)
-Arousal Work and Write (-0.10) Idle Tools (0.15)
-Productivity Idle Tools (-0.25) Social (0.20)
-Interruptions Lifestyle (-0.24) Communication Tools (0.42)
-Table 4. The computed features from smartphone usage behaviours
+All Valence Hybrid Work (-0.07) Tools Apps (0.11)
+Arousal Business Tasks (-0.05) Social Apps (0.15)
+Productivity Communication Apps (-0.24) Work-Related Communication (0.18)
+Interruptions Personal Tasks (-0.15) Communication Apps (0.29)
+Male Valence Communication Only(-0.08) Work-Related Communication (0.10)
+Arousal Device Idle (-0.12) Communication Only (0.11)
+Productivity Communication Apps (-0.15) Work-Related Communication (0.34)
+Interruptions Video Streaming (-0.23) Communication Apps (0.18)
+Female Valence Surfing (-0.17) Miscellaneous (0.13)
+Arousal Hybrid Work (-0.17) Social Apps (0.21)
+Productivity Communication Apps (-0.30) Personal Tasks (0.16)
+Interruptions Personal Tasks (-0.33) Device Idle (0.41)
+Age 18-24 Valence Surfing (-0.13) Tools Apps (0.09)
+Arousal Tools Apps (-0.17) Social Apps (0.19)
+Productivity Communication Apps (-0.21) Hybrid Work (0.20)
+Interruptions Social Apps (-0.10) Communication Apps (0.16)
+Age $\geq$25 Valence Communication Only (-0.12) Tools Apps (0.23)
+Arousal Hybrid Work (-0.10) Tools Apps (0.15)
+Productivity Tools Apps (-0.25) Social (0.20)
+Interruptions Personal Tasks (-0.24) Communication Apps (0.42)
+Interestingly, as shown in Table 4, we found mostly weak correlations, with a few activities exhibiting moderate relationships.
+For instance, activities related to Tools Apps (e.g., phone home screens and keyboards) showed a moderate positive correlation with valence, especially among older adults.
+This suggests that certain habitual or passive tasks, such as interacting with productivity tools, may positively influence mood, possibly linked to work-related or routine activities.
+Additionally, Surfing activity (e.g., browsing news, including COVID-19 updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood.
+The correlation with productivity was notably negative for Communication Apps (e.g., keyboards, chat apps), but positively correlated with Work-Related Communication activity (e.g., business apps like Microsoft Teams), highlighting that whether a communication tool is used for professional or personal purposes may influence productivity levels.
+Similarly, interruptions were mostly positively correlated with Communication Apps, particularly among older adults.
+This reflects the potentially disruptive nature of non-work-related communication during the day, which may vary by age.
+Conversely, Personal Tasks(e.g., personal apps or leisure activities) were negatively correlated with interruptions, especially for older adults and women, suggesting that non-work-related tasks may help buffer against disruptions, possibly offering a sense of control or relaxation.
+5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
+In this section, we address RQ2 ”Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”.
+First, we introduce the features extracted from mobile usage data and the prediction pipeline.
+We then list and discuss the prediction results of the general models and individual models.
+Finally, we examine the significant features for predicting human well-being.
+Table 5.
+The computed features from smartphone usage behaviours.
+For the valence model, the features selected by F-statistic $\geq 5^{*}$ or $\geq 10^{**}$.
+For the arousal model, the features selected by F-statistic $\geq 5^{+}$ or $\geq 10^{++}$.
 Category Feature Description
@@
 time_period The time of the day (morning, afternoon, evening and midnight)
-sin_weekday Sine Transformation of weekday
+sin_weekday∗∗,+ Sine Transformation of weekday
 cos_weekday Cosine Transformation of weekday
-sin_time Sine Transformation of time period
-cos_time Cosine Transformation of time period
+sin_time∗∗,++ Sine Transformation of time period
+cos_time∗∗,++ Cosine Transformation of time period
 is_weekend Binary values describing whether is is weekend or not
-Notification num_noti Number of notifications
-num_noti_interacted Number of notifications interacted
-noti_cat Categorical Transformation of the number of notifications
+Notification num_noti∗∗,++ Number of notifications
+num_noti_interacted∗∗,++ Number of notifications interacted
+noti_cat++ Categorical Transformation of the number of notifications
 noti_intr_cat Categorical Transformation of the number of notifications interacted
-noti_interacted_pcg The percentage of notification interacted to the total number of notifications
+noti_interacted_pcg∗∗,++ The percentage of notification interacted to the total number of notifications
 num_communication Number of notifications from the Communication Application
-main_contact The main social relationship that contacted with
+main_contact∗∗ The main social relationship that contacted with
 Physical Activity nunique_activity Number of unique physical activities
-main_activity The main physical activity of the users
-num_activity Number of physical activities
+main_activity∗∗,++ The main physical activity of the users
+num_activity∗,+ Number of physical activities
 App nunique_app The number of unique apps
-num_app The total number of apps
-nunique_app_cat The number of unique app categories
-app_noti The number of notification generated by non-communication apps
-nunique_app_noti The number of unique apps sent notifications
-main_cat_app The main category of apps
-second_cat_app The second main category of apps
+num_app∗∗,++ The total number of apps
+nunique_app_cat∗∗ The number of unique app categories
+app_noti∗∗,++ The number of notification generated by non-communication apps
+nunique_app_noti∗∗,++ The number of unique apps sent notifications
+main_cat_app∗,++ The main category of apps
+second_cat_app∗∗,++ The second main category of apps
 Activity Inferred from Apps nunique_user_activity The number of unique user activities
@@
 entropy_user_activity The entropy of different categories of user activity
-6. Predicting User Emotion and Social Roles from Mobile Usage Data
-In this section, we address RQ3 ”Can we model people’s emotion from mobile usage during home confinement?
-If so, what are the most significant types of features for prediction?”.
-First, we introduce the features extracted from mobile usage data and the prediction pipeline.
-We then list and discuss the prediction results of the general models and individual models.
-Finally, we examine the significant features for predicting human well-being.
 5.1. Feature Extraction
 Table 5 provides an overview of the features utilized in our research.
-Similar to previous mobile sensing studies (50; 81; 96), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
+Similar to previous mobile sensing studies (40; 69; 86), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
 To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences.
@@
 In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values.
-Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (44; 50), prompting us to conduct separate analyses for these two categories.
+Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (35; 40), prompting us to conduct separate analyses for these two categories.
 Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night.
@@
 Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
-To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 5.2.
+To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 4.3.2.
 AIGTS allows us to extract both activities and their corresponding durations.
@@
 5.2. Prediction Pipeline
-while emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (70; 98).
+While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (58; 88).
 We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance.
 In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
-Emotion.
+Emotion
 The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study.
 Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’.
-Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 means ‘excited’.
+Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 meaning ‘excited’.
 To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
-Regressors.
-In this research, we aim to develop regression models that accurately predicts human emotions during home confinement.
-To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (84) regressor, Random Forest (RF) (83) regressor, and Gradient Boosting (GB) (29) regressor.
+Regressors
+In this project, we aim to develop regression models that accurately predict human emotions during home confinement.
+To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (72), Random Forest (RF) (71), and Gradient Boosting (GB) (21).
 KNN is a widely utilized model known for its versatility across various domains.
@@
 GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
-Feature Selection.
+To explore predictive capacity more comprehensively, we first develop general models using data from all participants combined.
+Next, recognizing that participants with similar routines or lifestyles may exhibit comparable patterns, we construct group models by dividing participants based on their self-reported habits.
+Finally, we analyze individual models to account for the unique characteristics and patterns exhibited by different individuals.
+Feature Selection
 Various methods, such as SHAP (44), LIME (65), and the $F$-statistic ($F$) (60), can be used for feature selection.
@@
 Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable.
-In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and features selected based on an $F$ value greater than or equal to 10.
-Validation.
-To compensate for the limited richness of our data, we have implemented nested cross-validation (CV) (61) to train and evaluate our prediction model.
-This approach allows us to accurately estimate the unbiased generalization performance of the model.
-The nested CV consists of an inner loop and an outer loop.
-Initially, the dataset is divided into training, validation, and testing sets.
-This helps optimize the model’s performance.
-In the outer loop, similar to previous human-centred studies (33; 25), we divide the data into multiple groups, each group representing one participant.
-To ensure unbiased evaluation, we apply k-fold cross validation (93) where k=5 on all groups.
-This means that the data from the same participants are not used in both the training and test sets simultaneously.
-In the inner loop, the remaining data groups are further split into three folds, with each fold serving as a validation set.
-We optimize the hyperparameters on the training set, evaluate them on the validation set, and select the best parameter settings based on the results from the three folds.
-(a) MAE for valence
-(b) MAE for arousal
-(c) RMSE for valence
-(d) RMSE for arousal
-Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents standard deviation.
-Baselines.
-Similar to previous human-centered studies (33; 92), we compare the proposed model with two baselines.
-The first baseline is the Mean baseline, which calculates the average score of user responses.
-The second baseline is the Median baseline, which calculates the median value based on the data distribution and treats it as a predicted value.
+In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and greater than or equal to 10.
+Validation
+To address the limited richness of our data, we implemented Nested Cross-Validation (CV) (50) to train and evaluate our prediction models.
+Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation (83) with different values of $k$.
+The outer loop splits the data into $k$ folds, using each fold for testing.
+The inner loop further divides the remaining data into $k$ folds, with each fold serving as a validation set to train the model and optimize hyperparameters.
+The best parameter settings are selected based on the inner loop results.
+This approach provides an accurate estimate of the model’s unbiased generalization performance.
+For the general models, we use a 3-fold CV in the inner loop and a 5-fold CV in the outer loop, repeated over 3 trials.
+The group models employ 2-fold CV in the inner loop and 3-fold CV in the outer loop.
+Given the limited data available for individual participants, the individual models use a simpler 5-fold non-nested CV.
+Baselines and Metrics.
+Similar to previous human-centered studies (24; 82), we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value.
 To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (13).
@@
 5.3. Prediction Result
-To address RQ3, we proposed several regression models in the experiment.
-First, a general prediction model was constructed by considering all participants in the study.
-Second, we built individual prediction models for each participant to account for the unique characteristics and patterns exhibited by different individuals.
-It is worth mentioning that only mobile phone data was utilized in this study.
-After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis.
 5.3.1. General Model
-The prediction performance for the general models is presented in Figure 8.
-The models used in the comparison include Mean baseline, Median baseline, KNN regressor, RF regressor, GB regressor.
-Among these models, the RF regressor demonstrated superior performance to both baselines in terms of MAE and RMSE for predicting valance and arousal.
-Interestingly, the RF regressors with different feature sets achieved comparable performance, with all feature sets yielding the best results.
-This can be attributed to the capability of RF to assign weights to features based on their importance, where incorporating more features leads to improved prediction performance.
-Furthermore, it is observed that the proposed model generally exhibited better performance in predicting valence compared to arousal, the difference in performance may be attributed to the complexity of arousal, which is influenced by various factors such as physiological arousal and cognitive processing.
-These factors may not be fully captured by mobile usage behaviours alone.
+After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis were used to build the general models.
 Table 6.
@@
 […]
-Table 5 displays the prediction results of the general models for valence and arousal.
+The prediction performance for the general models is presented in Table 6.
+The models used in the comparison include Mean baseline, Median baseline, KNN regressor, RF regressor, GB regressor.
+Among these models, the RF regressor demonstrated superior performance to both baselines in terms of MAE and RMSE for predicting valance and arousal.
+Interestingly, the RF regressors with different feature sets achieved comparable performance, with all feature sets yielding the best results.
+This can be attributed to the capability of RF to assign weights to features based on their importance, where incorporating more features leads to improved prediction performance.
+Furthermore, it is observed that the proposed model generally exhibited better performance in predicting valence compared to arousal, the difference in performance may be attributed to the complexity of arousal, which is influenced by various factors such as physiological arousal and cognitive processing.
+These factors may not be fully captured by mobile usage behaviours alone.
 Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets.
@@
 These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
+5.3.2. Group Models
+We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: $EARLY$ Workers (678 data points from 12 participants who typically commence work between 7:00 and 10:00 AM) and $LATE$ Workers (778 data points from 13 participants who begin work after 3:00 PM).
+For the early workers, the average self-reported $valence$ is 3.2 (±$\pm 0.9$) and $arousal$ is 2.9 (±$\pm 1.0$).
+For the late workers, the average $valence$ is 3.7 (±$\pm 1.0$) and $arousal$ is 2.6 (±$\pm 1.2$).
+Prediction models were subsequently developed for each group using all features.
+As shown in Figure 7(a), when predicting $valence$ for early workers, the models performed similarly to the baselines.
+However, for late workers, the RF regressor outperformed the Mean baseline, achieving a 13.82% reduction in RMSE and a 20.36% reduction in MAE.
+A similar pattern is observed in Figure 7(c).
+For $arousal$ prediction, the models demonstrated limited predictive ability for early workers but performed better for late workers, with the RF regressor reducing RMSE by 17.88% and MAE by 27.74% compared to the Mean baseline.
+Compared to the general models, there was no improvement in $valence$ prediction for early workers, whereas for late workers, the models showed slight improvements (5% in RMSE) in both $valence$ and $arousal$ predictions.
+(a) RMSE for valence
+(b) MAE for valence
+(c) RMSE for arousal
+(d) MAE for arousal
+Figure 7.
+Prediction performance of $valence$ and $arousal$ models for two user groups: $EARLY$ workers (typically starting work between 7:00 and 10:00 AM) and $LATE$ workers (starting work after 3:00 PM).
+The models are evaluated with 3-fold cross-validation, where the error bar presents the standard deviation.
+These results highlight the potential value of considering user groups and routines when studying emotions.
+Specifically, the findings suggest that early workers’ emotions may be influenced by factors beyond app usage, such as external environmental cues during the day.
+In contrast, late workers may exhibit a stronger link between app usage behaviours and emotions.
+Further insights will be explored in the later section (Section 5.4) by analyzing feature importance.
 5.3.3. Individual Models
 The individual models for both valence and arousal were trained using all 29 features.
-However, due to some participants having insufficient data recorded, the models were only trained on those with more than 50 data points.
-This resulted in a total of 13 participants and 1220 data points being used.
+Due to insufficient data for some participants, only those with over 50 data points were included, resulting in a total of 13 participants and 1,220 data points.
 The prediction results for these models, with the ID of participants indicated on the y-axis, can be seen in Figure 8.
@@
 (a) MAE for valence
-(b) MAE for arousal
-(c) RMSE for valence
+(b) RMSE for valence
+(c) MAE for arousal
 (d) RMSE for arousal
@@
 To interpret the key factors influencing the prediction of valence and arousal during home confinement, we examine the feature importance of the best-performing models, which in this case are the RF regressors.
-For brevity, we focus on the top 15 features for each model, as shown in Figure 10.
 The RF regressors utilize impurity-based feature importance, which measures how much including a particular feature reduces impurity across the decision tree nodes.
 A higher score indicates a more important feature, as it leads to a greater reduction in impurity.
-A higher score indicates a more important feature, as it leads to a greater reduction in impurity.
+For brevity, we focus on the top 15 features for the general model and compare them with the top 5 features for the group models.
+Figure 9. Top 15 features based on impurity score with RF regressors of the general models.
 As depicted in Figure 9, we observe that the most important features for both valence and arousal are similar, but with different weights.
-These features primarily include time-related features, app-category features, and notification features.
-These findings suggest that these variables play a significant role in accurately predicting participants’ valence and arousal levels during home confinement.
-(a) Valence
-(b) Arousal
-Figure 10. Top 15 features based on impurity score with RF regressors
-Both models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day significantly influences participants’ valence and arousal.
-Additionally, the day of the week (‘weekday’) impacts both models, although it has less significance in the arousal model.
-Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels.
+These features primarily include time-related, app-category, and notification features.
+Both general models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day greatly influences participants’ valence and arousal.
+Additionally, the day of the week (‘weekday’) impacts both models, although it has less importance in the arousal model.
+Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group models.
+It worth mentioning that ‘time_period’ is the most important feature for predicting $valence$ for $EARLY$ workers, but not for $LATE$ workers.
 These findings differ slightly from the observations reported in (35), which found that app usage patterns varied greatly by time and day.
@@
 This aligns with the findings presented in (12), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
-The above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement.
-They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods.
-By considering the impact of app usage, notifications, and time-related factors, we can also design targeted strategies to enhance individuals’ emotional well-being and overall quality of life.
-7. Implications
-Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance.
-Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles.
-First of all, our findings highlight the importance of mental health during periods of home confinement.
-Home confinement is not limited to temporary pandemic adaptation but also the growing trend of working from home post-pandemic (40; 69) and other similar situations, such as post-operative recovery, parental leave, or limited mobility related to disability, or even mobile phone addiction.
-Secondly, our quantitative approach and regression models can contribute to tracking, early detection and treatment of mental illnesses during these circumstances.
-Our results also help policymakers to develop appropriate strategies that minimize the long-term negative effects of the pandemic on mental health and encourage people to build healthy digital habits (69).
-For example, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al.
-(12).
-In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (53; 40).
-Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviours to those during the pandemic (57), underscoring the importance of studying people’s behaviours in high technology-reliance contexts.
-Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions.
-Furthermore, our study contributes to understanding the mental health benefits and consequences of working from home or relying heavily on technology.
-For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
-In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviours of individuals under home confinement conditions, which has important implications for various fields.
-For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation.
-Rethinking about the future of work and life scenarios, where heavily rely on digital technologies.
-It is important to understand how people deal with being confined to their homes and how technology impacts their lives.
-This understanding can help researchers design smart environments, or assist healthcare professionals in developing effective remote support programs that enhance the quality of care for those in similar situations.
-Furthermore, the findings from this study may have broader implications for our understanding of human behaviour and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
+Figure 10. Top 5 features based on impurity score with RF regressors of group models.
+Besides, as shown in Figure 10, for $LATE$ workers, the same five features with similar importance distributions are observed for both $valence$ and $arousal$ predictions, all are related to app interaction.
+In contrast, $EARLY$ workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contribute to $valence$ prediction.
+This difference may reflect the broader range of daytime activities available to $EARLY$ workers, potentially explaining their lower prediction performance discussed in Section 5.3.2.
+For $LATE$ workers, the reliance on app interaction highlights a potential concern regarding phone dependency.
+6. Discussion and Implications
+This study explores the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors.
+Although the study specifically focuses on the COVID-19 lockdown, we argue that its findings have broader relevance, especially in light of the growing trend of remote work and other situations of enforced isolation.
+6.1. Well-being and Social Roles During Home Confinement (RQ1)
+Our analysis revealed that participants generally experienced a state of calmness and relaxation, reflected in relatively high levels of valence (positive emotional state) and lower levels of arousal (indicating a lack of intense emotional stimulation).
+While this suggests a degree of emotional stability during the lockdown, the study also highlights the negative effects of home confinement on work-life balance.
+Participants reported longer work hours and the blurring of personal and professional boundaries, which resonates with the challenges faced by remote workers in post-pandemic contexts.
+The segmentation of app usage data also identified key behaviors, such as increased use of communication apps and idle device time, which may reflect both increased virtual socialization and available free time.
+Although the findings are rooted in the context of the COVID-19 lockdown, the implications extend beyond this unique situation.
+Home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction.
+The growing trend of telework and flexible work arrangements in the post-pandemic world further underscores the importance of understanding the emotional and behavioral impacts of prolonged confinement in various forms.
+6.2. Predicting Emotional States Through Mobile Usage (RQ2)
+In response to RQ2, we developed regression models that predict participants’ valence and arousal based on their mobile usage data.
+These models demonstrated strong predictive performance, emphasizing the value of mobile phones as a tool for tracking emotional well-being in real time.
+Our findings indicate that time-related features (e.g., the timing of mobile phone usage) were key predictors of both valence and arousal, followed by app notifications and interactions with those notifications.
+This temporal sensitivity suggests that personalized interventions based on users’ specific routines could enhance emotional monitoring and support during home confinement.
+One key finding was the variability in the performance of prediction models across different user groups.
+For example, users who start work later in the day relied more heavily on mobile phone interactions, while those with early start times would benefit from models that incorporate data from other activities, such as physical exercise or interactions with environmental cues.
+This insight is crucial in extending the applicability of our findings to various forms of isolation beyond the pandemic.
+Whether it’s remote workers, individuals recovering from surgery, or people experiencing limited mobility due to health conditions, the need for individualized and context-aware models remains a central takeaway.
+6.3. Design Implications
+Based on the findings of this study, we propose several design implications for future interventions and mobile applications aimed at supporting well-being during periods of home confinement or isolation:
+Personalized Predictive Models.
+The success of time-related features in predicting emotional states suggests that mobile apps should incorporate personalized, time-sensitive data to offer real-time emotional tracking and tailored interventions.
+For example, apps could adjust their notifications and suggestions based on users’ routines, ensuring that interventions are timely and relevant.
+Balanced Digital Interactions.
+Our study found that interruptions, particularly from communication apps, may exacerbate the blurring of work-life boundaries.
+Future apps should focus on helping users manage interruptions, offering customizable notification settings or “focus mode” features to reduce disruptions and help maintain better separation between work and personal life.
+Encouraging Healthy Digital Habits.
+Tools that support productivity, such as task managers and calendars, were associated with higher valence.
+Future designs should encourage healthy digital habits by integrating features that promote structure, goal-setting, and task management.
+Additionally, incorporating reminders for physical activity or mental health breaks could further support users’ emotional well-being during periods of isolation.
+Context-Aware Interventions.
+Given the differences in how users experience isolation based on their routines, digital interventions should be context-aware and adaptive.
+For example, users with later work hours may benefit from interventions encouraging physical activity or relaxation earlier in the day, while users with earlier start times may need more focused strategies for managing digital distractions during the day.
+6.4. Generalizability of the Findings
+Although this study specifically focuses on the COVID-19 home confinement context, we believe the findings can extend to a range of other isolation scenarios.
+While the emotional triggers—such as fear of infection in a pandemic or frustration with limited mobility in recovery—may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar.
+Our study observed patterns like emotional detachment, stress, and blurred work-life boundaries, all of which are commonly experienced during prolonged isolation, regardless of the specific circumstances.
+The emotional dynamics of isolation—such as changes in social interactions, work-related stress, and adjusting to new routines—are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement.
+For example, both remote workers and those recovering from surgery face challenges with work-life balance and social isolation.
+Our findings about app usage—especially productivity apps promoting positive emotional states—are relevant in these contexts, suggesting that digital tools can support well-being across various isolation scenarios.
+Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond COVID-19.
+While the types of apps and behaviors may vary, the principle of using mobile devices to track emotional states in real-time is adaptable to different situations, such as remote work or recovery periods.
+The ability to monitor and intervene based on individual usage patterns could be valuable in a wide range of confinement contexts.
+Lastly, context-aware interventions are crucial for emotional well-being during isolation.
+Our study highlights the importance of personalized, real-time interventions—whether it’s through managing digital interruptions, promoting relaxation, or encouraging physical activity.
+These strategies can be adapted for any form of isolation, from remote work to recovery, offering broader applicability for improving mental health.
 7. Limitations
-We have identified three main limitations in this research.
-One significant limitation is the selection of time and place.
-Regarding time, we were unable to compare user well-being and digital device usage data before COVID-19 lockdown with the data during the lockdown due to the unpredictable nature of lockdown policies.
-However, we have mitigated this limitation by allowing users to report the impact of the lockdown on their well-being and work-life balance through self-reported questionnaires.
-These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4.
-In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage IV restrictions was applied.
-While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (11; 15; 12).
-Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
-The second limitation is the relatively short duration of the data collection period.
-We chose to collect data for only three weeks in order to minimize interruptions from ESM questionnaires and end-of-day surveys, with the goal of reducing the burden on participants during the challenging realities of COVID-19 home confinement.
-Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses.
-It is worth noting that the delivery of the end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience.
-Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data.
-However, due to the limited three-week period of our data collection, it is important to acknowledge that if strict home confinement lasts longer, the emotional experiences of individuals may evolve as they adapt to new challenges.
-Finally, the third limitation of this study is the number and representativeness of the participants.
-A total of 32 participants were recruited, with 27 using smartphones, 13 using desktops, and only 8 using both.
-Since the majority of participants use mobile phones for data collection, the majority of our background application data analysis focuses on mobile usage data.
-Additionally, participants were primarily recruited through Facebook and Discord advertisements, which may have led to a population that is skewed towards tech-savvy information workers.
-Additionally, there is a possibility of oversampling individuals who were ‘relieved to be safe at home and cherish the opportunity to spend more quality time with their family,’ as these individuals may have had more availability to participate in the study.
-It is important to acknowledge that the COVID-19 lockdown presented challenges in recruiting participants.
-However, despite these challenges, we were able to collect sufficient data from each participant to conduct our analysis.
+This study has several limitations that must be acknowledged, which may affect the validity of our findings (78):
+Time and Geographical Constraints (Threats to Construct and External Validity).
+The timing and location of data collection limit the generalizability of our results.
+We could not compare users’ well-being and digital device usage before and during the COVID-19 lockdown due to the unpredictability of lockdown policies.
+Although participants reported their experiences via self-reported questionnaires, the effects of different lockdown measures and their duration may vary across regions.
+As the study was limited to Melbourne, the findings may not fully reflect experiences in other areas with differing restrictions.
+Short Duration of Data Collection (Threats to Internal Validity).
+The data collection period was limited to three weeks to minimize participant burden during home confinement.
+While this yielded a substantial dataset, a longer observation period would have provided deeper insights into how digital device usage and emotional well-being evolve over time, particularly as individuals adapt to prolonged lockdowns.
+Future studies should consider longer data collection windows to capture these potential changes.
+Device Usage and Untracked Data (Threats to Internal and Construct Validity).
+A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops while only contributing smartphone data.
+These unmonitored interactions could have influenced emotional states and introduced noise into the data.
+The absence of comprehensive tracking across all devices limits our ability to fully assess participants’ total digital behavior and emotional responses, suggesting the need for more extensive device tracking in future studies.
+Sample Representativeness (Threats to External and Construct Validity).
+The sample size (32 participants) and recruitment methods (via Facebook and Discord ads) may have resulted in a skewed population, likely favoring tech-savvy individuals, particularly information workers.
+Additionally, participants who felt more positively about being at home during the lockdown may have been overrepresented.
+These factors limit the generalizability of the findings to a broader population.
+However, despite these challenges, the data collected was sufficient for analysis.
 8. Conclusion
-In this study, we delved into how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period.
-Our proposed models yielded promising results in predicting human emotions during home confinement.
-Moreover, we uncovered intriguing insights regarding the correlation between different factors and human well-being.
-To address our research objectives on home confinement, we formulated three research questions and key findings are summarized below.
-In response to RQ1, we found that participants reported relatively high levels of valence but relatively low levels of arousal, suggesting that participants experienced a state of calmness and relaxation during COVID-19 home confinement.
-Additionally, our analysis suggested that the COVID-19 lockdown has a detrimental effect on valence and work-life balance, with participants experiencing longer work hours due to interruptions and a blurring of personal and work boundaries.
-Regarding RQ2, we segmented app usage data and identified the most frequent user activities ‘Just Communication’ and ‘Idling’, potentially attributed to increased virtual communication and available free time.
-We also found that tools-related activities positively correlated with valence, suggesting people used productivity apps effectively.
-Moreover, interruptions displayed a positive association with Communication Tools, possibly due to the social nature of these tools leading to more interruptions during work hours.
-Moving on to RQ3, we developed regression models to predict people’s valence and arousal based on their mobile usage during home confinement.
-These models exhibited superior performance compared to baseline models.
-Notably, our findings revealed that time-related features emerged as the primary predictors for both valence and arousal.
+In this field study, we explored how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period.
+This rare event offers optimal conditions to while minimizing confounding factors like nature and physical social interactions.
+It contributes to understanding the mental health implications of heavy reliance on technology and work-from-home arrangements.
+By tracking mental well-being, the study highlights how individuals manage work-life balance and adapt their schedules or arrangements (e.g., work-from-home or half-office half-home) effectively.
+Additionally, it provides valuable insights into psychological states and smartphone usage behaviours during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation.
+Understanding how people navigate home confinement and the role of technology in their lives can guide the design of smart environments and remote support programs, improving care quality.
+These findings also have broader implications for understanding human behaviour during crises, supporting the development of policies and interventions for future challenges.
 Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviours during home confinement is crucial.

### v4

@@
-”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement
-Abstract.
+Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement
+Abstract
 Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life.
-While essential in various scenarios, confinement within the home environment can profoundly influence mental well-being and digital device usage.
-Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of prolonged home confinement on mobile device usage.
+While essential in various scenarios, it can profoundly influence mental well-being and digital device usage.
+Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of home confinement on mobile device usage.
 We conducted an in-situ study with 32 participants, analyzing three weeks of mobile usage data to assess emotional well-being and social dynamics in restricted environments.
-Our findings reveal that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation.
-This study highlights the potential for developing targeted interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
-Emotion, Home Confinement, Mobile Sensing, COVID-19 Lockdown, Social Roles, Well-Being
-copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
-1. Introduction
-Home confinement, defined as the restriction of movements within the home (52) due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives (38; 11).
-Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression (6; 66).
+We found that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation.
+This study highlights the potential for developing interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
+keywords:
+Human-centred Computing , Mobile Sensing , Mental wellbeing , Affective computing , Home confinement , Emotion
+journal: Computers in Human Behavior
+[…]
+[…]
+[…]
+[aff3]organization=RMIT University, city=Melbourne, postcode=3000, country=Australia
+[…]
+[aff_g]organization=University of Kassel, city=Kassel, postcode=34131, country=Germany
+1 Introduction
+Home confinement, defined as the restriction of movements within the home [1] due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives [2, 3].
+Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression [4, 5].
 For instance, Bartoszek et al.
-(6) found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue.
+[4] found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue.
 Understanding these impacts is essential for developing effective targeted interventions to support individuals during such periods.
-Traditionally, mental well-being during home confinement has been assessed using survey-based instruments (66; 22), which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias (23; 59).
-Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative (61; 15).
-While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation (61; 15), they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited (61).
+Traditionally, mental well-being during home confinement has been assessed using survey-based instruments [5, 6], which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias [7, 8].
+Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative [9, 10].
+While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation [9, 10], they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited [9].
 This study shifts focus to more granular data, such as app usage patterns, to provide deeper insights into emotional well-being during confinement.
@@
 Factors such as the reason for confinement, personal health conditions, and social environment can vary greatly, complicating the creation of a standardized model.
-However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies (57; 70).
+However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies [11, 12].
 This shared experience allows for a more uniform analysis of how mobile device usage reflects emotional states, social roles, and work-life dynamics during such periods.
-Espeically, while this study uses the lockdown as a descriptive case study (87), the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
-In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles, during COVID-19 lockdown, as a case study for home confinement.
-We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health (17; 54).
-By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect (67) and Social Role from the Role Theory (8), we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics.
+Especially, while this study uses the lockdown as a descriptive case study [13], the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
+In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles during home confinement, using data collected during the COVID-19 lockdown as a case study.
+We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health [14, 15].
+By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect [16] and Social Role from the Role Theory [17], we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics.
 Our research seeks to address the following questions: RQ1.
-How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?
+How are emotions and social roles during home confinement perceived through self-reports and inferred from app usage patterns?
 RQ2.
@@
 To answer these questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia, involving 32 participants.
-We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) (7) and end-of-day surveys.
+We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) [18] and end-of-day surveys.
 We also proposed a novel method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records by identifying segments and transition times between apps during the day.
 The resulting insights offer a deeper understanding of how app usage during confinement reflects emotional well-being and social dynamics, with implications for developing digital interventions to support individuals in long-term home confinement.
-All scripts for this study are open source.
-[^1] Link will be shared after the double-blind review.
-In summary, our contributions are as follows:
+All scripts for this study are open source. [^4] Link will be shared after the double-blind review. In summary, our contributions are as follows:
+  • 1.
     We collected a rich dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia.
-This dataset includes 502,485 records of users smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys.
+This dataset includes 502,485 records of users’ smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys.
 It provides a unique, real-world snapshot of emotional well-being and social dynamics during lockdown.
+  • 2.
     We analyzed perceived valence, arousal, and social roles, revealing that the lockdown significantly impacted well-being, reducing emotional valence and increasing work hours due to interruptions or blurred work-life boundaries.
 These findings shed light on the complex emotional landscape during prolonged confinement.
+  • 3.
     We introduced the App-based Information Gain Temporal Segmentation (AIGTS) method, an unsupervised technique for extracting user activities (e.g., video-watching, socializing, deep work) from app usage data.
-This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviours relate to emotional well-being.
+This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviors relate to emotional well-being.
+  • 4.
     We demonstrated that mobile usage patterns can predict emotional well-being during home confinement.
 Our extensive experiments revealed key features that are most predictive of emotional states, offering new avenues for developing digital interventions to support individuals during isolation.
-2. Related work
-2.1. Impact of the Home Confinement on Mental Well-Being
-Home confinement, whether due to edical conditions, disability, or societal events like pandemics, can significantly affect mental well-being.
-Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness (76; 14; 10).
-The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger (64; 39).
-These effects may persist even after restrictions are lifted (31).
+2 Related work
+2.1 Impact of the Home Confinement on Mental Well-Being
+Home confinement, whether due to medical conditions, disability, or societal events like pandemics, can significantly affect mental well-being.
+Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness [19, 20, 21].
+The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger [22, 23].
+These effects may persist even after restrictions are lifted [24].
 The COVID-19 pandemic underscored these effects on a global scale.
-The World Health Organization (84) reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks.
-Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread (85; 16; 37).
+The World Health Organization [25] reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks.
+Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread [26, 27, 28].
 For example, Khubchandani et al.
-(36) found depression (39%) and anxiety (42%) surged in the U.S.
+[29] found depression (39%) and anxiety (42%) surged in the U.S.
 post-lockdown.
 Similarly, Fountoulakis et al.
-(20) reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece.
-Loneliness was also a major issue, especially among those in long-term isolation (37).
+[30] reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece.
+Loneliness was also a major issue, especially among those in long-term isolation [28].
 Despite extensive research on the mental well-being impacts of various forms of home confinement, much of it is geographically limited or reflects different phases of isolation.
 This variability makes it challenging to fully understand the broader, long-term psychological effects of home confinement in diverse contexts.
-To address this gap, RQ1 examines how emotions and social roles are experienced during COVID-19 home confinement, using self-reports and app usage patterns to infer these dynamics.
-2.2. Digital Device Usage During Home Confinement
+To address this gap, RQ1 examines how emotions and social roles are experienced during home confinement, using self-reports and app usage patterns to infer these dynamics.
+2.2 Digital Device Usage During Home Confinement
 During periods of home confinement, digital devices play a crucial role in maintaining communication, accessing information, and fulfilling work and social roles.
-With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities (32).
+With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities [31].
 For instance, Ratan et al.
-(62) found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
-While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behaviour under prolonged confinement.
-With workplaces and schools shifting to remote environments, the demand for digital communication tools surged (79; 19; 5).
-Video conferencing platforms, social media, and messaging apps saw massive increases in usage (70; 57), and overall digital device use skyrocketed.
-A global survey found 70% of internet users reported increased smartphone usage during the lockdown (70), emphasizing how integral smartphones became in managing both work and personal life.
+[32] found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
+While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behavior under prolonged confinement.
+With workplaces and schools shifting to remote environments, the demand for digital communication tools surged [33, 34, 35].
+Video conferencing platforms, social media, and messaging apps saw massive increases in usage [12, 11], and overall digital device use skyrocketed.
+A global survey found 70% of internet users reported increased smartphone usage during the lockdown [12], emphasizing how integral smartphones became in managing both work and personal life.
 Similarly, Majumdar et al.
-(45) observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown.
+[36] observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown.
 Jonnatan et al.
-(33) also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
-In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication (43; 29).
-Studies suggest that behaviours such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase (46).
+[37] also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
+In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication [38, 39].
+Studies suggest that behaviors such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase [40].
 This highlights the importance of understanding digital device usage patterns in today’s technology-reliant environments, particularly in contexts of home confinement.
-2.3. Inferring Mental Well-Being Using Mobile Sensing Technologies
-Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood (49; 88; 48), engagement (24; 28; 2), and mental health issues like depression and anxiety (15; 61; 81).
+2.3 Inferring Mental Well-Being Using Mobile Sensing Technologies
+Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood [41, 42, 43], engagement [44, 45, 46], and mental health issues like depression and anxiety [10, 9, 47].
 Many studies have integrated data from multiple sensors to improve prediction accuracy.
 For example, Gao et al.
-(24) used wearable and environmental sensors to predict student engagement, while Wang et al.
-(81) tracked depression using passive mobile sensing and wearables.
+[44] used wearable and environmental sensors to predict student engagement, while Wang et al.
+[47] tracked depression using passive mobile sensing and wearables.
 Jaques et al.
-(30) modelled student mood using physiological signals and mobile behaviour.
-While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviours like physical activity, communication patterns, and location (25).
-Studies have shown that mobile usage data can predict psychological states such as anxiety (61; 77), mood (88; 48), and emotion (75).
+[48] modeled student mood using physiological signals and mobile behavior.
+While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviors like physical activity, communication patterns, and location [49].
+Studies have shown that mobile usage data can predict psychological states such as anxiety [9, 50], mood [42, 43], and emotion [51].
 For instance, Zhang et al.
-(88) developed an app to detect compound emotions from smartphone usage, and Tlachac et al.
-(77) used call and text logs to predict anxiety and depression during the pandemic.
-Research on app sequences has also provided insights into daily behaviours, such as social roles (3) and interruptibility (55; 56).
+[42] developed an app to detect compound emotions from smartphone usage, and Tlachac et al.
+[50] used call and text logs to predict anxiety and depression during the pandemic.
+Research on app sequences has also provided insights into daily behaviors, such as social roles [52] and interruptibility [53, 54].
 Anderson et al.
-(3) demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al.
-(55; 56) used app sequences to predict activity breakpoints for intelligent notification delivery.
+[52] demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al.
+[53, 54] used app sequences to predict activity breakpoints for intelligent notification delivery.
 However, most studies have focused on identifying breakpoints rather than modeling user activity or predicting contextual variables.
@@
 While mobile sensing technologies have shown potential in predicting well-being, gaps remain in understanding how these tools can be applied specifically during home confinement.
-Home confinement introduces unique behaviours—such as increased reliance on digital devices and altered perceptions of time—that differ from daily life.
-Although some studies (77; 53; 74) have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement.
+Home confinement introduces unique behaviors, such as increased reliance on digital devices and altered perceptions of time, that differ from daily life.
+Although some studies [50, 55, 56] have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement.
 Understanding this unique setting could provide valuable insights for other confinement scenarios, such as post-operative recovery, parental leave, or aging.
 Therefore, RQ2 investigates how mobile sensing can model emotions during home confinement and identify key features for prediction.
-3. Study Overview
+3 Study Overview
 We conducted a three-week field study in Melbourne, Australia, from 24 August 2020 to 13 September 2020, during the strictest Stage IV COVID-19 restrictions.
-This study use Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work.
+This study uses Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work.
 The following section details the background, participants, and data collected.
-3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia
-Figure 1. The periods of six lockdowns and our data collection
-To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (47; 73).
+3.1 A Case Study of COVID-19 Lockdown in Melbourne, Australia
+Figure 1: The periods of six lockdowns and our data collection
+To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world [57, 58].
 Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1).
-Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (51).
+Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne [59].
 Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons.
@@
 On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions.
-During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (62; 70).
-3.2. Procedures
+During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication [32, 12].
+3.2 Procedures
 Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers.
@@
 However, as our study focused on psychological states and digital device usage during COVID-19 home confinement, not social media behaviour, we believe this does not significantly impact the validity of our findings.
-Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps(27; 80) on their smartphones and desktops.
+Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps[60, 61] on their smartphones and desktops.
 These apps, designed to record usage behaviors and enable experience sampling, were optimized for minimal computational, energy, and storage consumption.
@@
 Participants were informed of their right to withdraw at any time without penalty, and their data would be excluded from the study upon request.
-Table 1. Overview of the collected data records across devices
+Table 1: Overview of the collected data records across devices
 Category Device Total Records Participants Data Used for RQs
-App Both 502,485 32 RQ1
+App & Both 502,485 32 RQ1
 Desktop 261,911 13 /
 Smartphone 240,574 27 RQ2
-ESM Both 1,749 28 RQ1
+ESM & Both 1,749 28 RQ1
 Desktop 227 10 /
@@
 Ene-of-Day Survey Both 265 25 RQ1
-Table 2. Detailed overview of data categories, items, and response options
+Table 2: Detailed overview of data categories, items, and response options
 Category Items Questions / Descriptions Options
@@
 Difference How was today different from yesterday? Text (optional)
-3.3. Collected Data
+3.3 Collected Data
 The data collection period spanned 21 days.
@@
 ESM responses.
-There are some popular questionnaires to measure well-being (34; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research.
+There are some popular questionnaires to measure well-being [62, 63], and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) [63] was employed in this research.
 In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and tasks from users (see Table 2).
@@
 and rated with a 5-point Likert scale, from unhappy to happy for valence and from calm to excited for arousal.
-These measures were used to represent the users’ emotions based on the circumplex model (67).
-The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework (4).
+These measures were used to represent the users’ emotions based on the circumplex model [16].
+The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework [64].
 (a) ESM
 (b) End-of-day survey
-Figure 2. The screenshots for the ESM and End-of-day survey
+Figure 2: The screenshots for the ESM and End-of-day survey
 End-of-day survey.
@@
 Items 3, 4, 5, and 9 were rated from 1 to 5, indicating a great deal, a lot, a moderate amount, a little, and not at all.
-4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement (RQ1)
-In this section, we address RQ1 ”How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?”.
+4 Understanding Emotions, Social Roles, and Work-life-balance During Home Confinement (RQ1)
+In this section, we address RQ1 “How are emotions and social roles during home confinement perceived through self-reports and inferred from app usage patterns?”.
 First, we analyse 1522 ESM responses collected via smartphones from 25 participants.
-Next, we explore the impact of the COVID-19 lockdown by examining the end-of-day survey.
+Next, we explore the impact of the lockdown by examining the end-of-day survey.
 Additionally, we explore users’ app usage behaviours through unsupervised activity recognition, identifying common activity clusters and their relationship to the perceived psychological metrics.
-4.1. The Perceived Valence, Arousal and Social Roles During Covid-19 Home Confinement
+4.1 The Perceived Valence, Arousal and Social Roles During Home Confinement
 Figure 3 illustrates the distribution of perceived valence, arousal, and social roles.
@@
 Furthermore, the solidarity and support networks that emerged during the pandemic likely provided additional emotional comfort.
-However, it is important to note that our data collection occurred during a three-week period as the COVID-19 Stage IV restrictions were lifting.Had the confinement continued longer, participants likely would have reported more negative emotions, such as boredom or frustration, as they adjusted to the new normal.
-Next, we investigate the role of social roles in shaping emotional experiences during home confinement.
-A social role refers to the expectations, norms, and behaviours associated with particular social contexts (63).
-These roles usually influence users’ behaviours and the type of information they prefer to receive (4).
-As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
-We further explore how emotions vary across different social roles.
-Figure 3(c) shows that participants reported the highest valence and lowest arousal when engaged in private-related roles, compared to work-related or hybrid roles.
-This suggests that personal activities, which are often less stressful and more enjoyable, foster a greater sense of well-being.
-Conversely, work-related and hybrid roles were associated with higher levels of stress and pressure.
-Welch’s t-tests (89) confirm these findings, showing significant differences in valence between private-related and the other two roles (i.e., private vs.
-both, private vs.
-work), with p-value ¡ 0.05.
-The findings offer insights into how individuals can structure their daily routines to boost positive emotions during home confinement.
-Interventions aimed at improving well-being in such situations may benefit from considering the social roles individuals are engaging in (26).
+However, it is important to note that our data collection occurred during a three-week period as the COVID-19 Stage IV restrictions were lifting.
+Had the confinement continued longer, participants likely would have reported more negative emotions, such as boredom or frustration, as they adjusted to the new normal.
 (a) Valence and arousal
@@
 (c) Social roles and emotion
-Figure 3. The distribution of perceived valence, arousal, social roles and their relationship
-4.2. The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
+Figure 3: The distribution of perceived valence, arousal, social roles and their relationship
+Next, we investigate the role of social roles in shaping emotional experiences during home confinement.
+A social role refers to the expectations, norms, and behaviours associated with particular social contexts [65].
+These roles usually influence users’ behaviours and the type of information they prefer to receive [64].
+As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
+[…]
+4.2 The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
 End-of-day survey responses reveal that the lockdown had a noticeable impact on participants’ valence, arousal, and work-life balance.
@@
 Common complaints included missing social interactions (9 responses), the inability to go outside (10 responses), and a sense of boredom (6 responses).
-For instance, ”I thought I might be able to see friends, but I did not”, ”I want to go outside, to the park”, ”I’m missing my family”, ”It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long.
+For instance, “I thought I might be able to see friends, but I did not”, “I want to go outside, to the park”, “I’m missing my family”, “It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long.
 I’m bored and sick of my housemates and it feels like time has stopped”.
 On the other hand, two participants reported positive effects, such as feeling happier or using the extra time for physical activity and seamless work transitions.
+Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., “I don’t feel motivated to do productive things”, “Emotionally tired, perhaps as a result of the lockdown”.
+However, two participants noted that they felt more energized during the lockdown.
+In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as “I worked a bit more than I should’ve”, “Everything blends together, separating work and life is really difficult”.
+These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities.
+While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
 (a) Productivity
@@
 (c) Influence on well-being
-Figure 4. The distribution of productivity, interruption, and influence during the lockdown
-Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., ”I don’t feel motivated to do productive things”, ”Emotionally tired, perhaps as a result of the lockdown”.
-However, two participants noted that they felt more energized during the lockdown.
-In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as ”I worked a bit more than I should’ve”, ”Everything blends together, separating work and life is really difficult”.
-These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities.
-While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
-4.3. Understanding App Usage Behaviours During COVID-19 Home Confinement
-Table 3. App categories (referred to the Google Play Store categories) and top three frequently used apps.
+Figure 4: The distribution of productivity, interruption, and influence during the lockdown
+4.3 Understanding App Usage Behaviours During Home Confinement
+Table 3: App categories (referred to the Google Play Store categories) and top three frequently used apps.
 Upper Category Subcategory Top App Second App Third App
@@
 Personalization One UI Home OnePlus Launcher TouchWiz home
-4.3.1. App Categorization
+4.3.1 App Categorization
 Apps were categorized based on historical usage logs.
@@
 These were grouped under the Tools category, along with utility apps like Explorer.
-4.3.2. Unsupervised Activity Recognition From App Sequences
+4.3.2 Unsupervised Activity Recognition From App Sequences
 To identify distinct user activities from app usage, we propose an unsupervised approach that combines segmentation and clustering.
@@
 Figure 5 illustrates how activities are extracted from app usage data.
-AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) (68), an unsupervised method for segmenting multivariate time series.
+AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) [68], an unsupervised method for segmenting multivariate time series.
 IGTS identifies splits that minimize the weighted entropy across the multivariate time series, where each variable corresponds to whether a specific app category was used.
@@
 These steps ensure that the average segment length remains consistent across all days.
-Figure 5. An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
+Figure 5: An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
 Clustering.
 Next, AIGTS is applied to the all-time series, resulting in a segmented list for each user.
-To extract features from these segments, K-means clustering (41) was applied.
+To extract features from these segments, K-means clustering [69] was applied.
 The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories.
-To identify the optimal number of segments, Silhouette Analysis (42) was utilized with a maximum of 20 clusters.
+To identify the optimal number of segments, Silhouette Analysis [70] was utilized with a maximum of 20 clusters.
 Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment.
@@
 In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
-Figure 6. Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
+Figure 6: Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
 Overview of User Activities.
@@
 The Device Idle cluster, corresponding to Personalization apps (e.g., launchers, as shown in Table 3), represents periods when users remain on their device’s home screen.
-We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the COVID-19 home confinement, accounting for 17% and 14.6% of all activities, respectively.
+We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the home confinement, accounting for 17% and 14.6% of all activities, respectively.
 These activities are characterized by high proportions of communication-related apps and personalization apps, suggesting users spent considerable time on their devices in passive states or communicating with others.
@@
 The differentiation between Light Work and Deep Work clusters highlights variations in the intensity and type of work-related app usage.
-4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
+4.3.3 Correlation Analysis of User Activity/App Usage and Well-Being
 .
-We further investigate how user activities and app usage behaviours correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions.
-To assess these correlations, we performed Spearman Rank Correlation (18) tests between the time spent on different activities/apps and the scores for each dimension.
-Specifically, valence and arousal were examined for theirr relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
-Table 4. Top correlations between app usage/user acitivty and well-being dimensions.
+We further investigate how user activities and app usage behaviors correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions.
+To assess these correlations, we performed Spearman Rank Correlation [71] tests between the time spent on different activities/apps and the scores for each dimension.
+Specifically, valence and arousal were examined for their relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
+Table 4: Top correlations between app usage/user activity and well-being dimensions.
 Groups Variable Lowest correlation $\rho$ Highest correlation $\rho$
@@
 This suggests that certain habitual or passive tasks, such as interacting with productivity tools, may positively influence mood, possibly linked to work-related or routine activities.
-Additionally, Surfing activity (e.g., browsing news, including COVID-19 updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood.
+Additionally, Surfing activity (e.g., browsing news, including pandemic updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood.
 The correlation with productivity was notably negative for Communication Apps (e.g., keyboards, chat apps), but positively correlated with Work-Related Communication activity (e.g., business apps like Microsoft Teams), highlighting that whether a communication tool is used for professional or personal purposes may influence productivity levels.
@@
 Conversely, Personal Tasks(e.g., personal apps or leisure activities) were negatively correlated with interruptions, especially for older adults and women, suggesting that non-work-related tasks may help buffer against disruptions, possibly offering a sense of control or relaxation.
-5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
-In this section, we address RQ2 ”Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”.
+5 Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
+In this section, we address RQ2 “Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”.
 First, we introduce the features extracted from mobile usage data and the prediction pipeline.
@@
 Finally, we examine the significant features for predicting human well-being.
-Table 5.
-The computed features from smartphone usage behaviours.
+Table 5: The computed features from smartphone usage behaviours.
 For the valence model, the features selected by F-statistic $\geq 5^{*}$ or $\geq 10^{**}$.
@@
 entropy_user_activity The entropy of different categories of user activity
-5.1. Feature Extraction
+5.1 Feature Extraction
 Table 5 provides an overview of the features utilized in our research.
-Similar to previous mobile sensing studies (40; 69; 86), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
-To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences.
+Similar to previous mobile sensing studies [72, 73, 74], we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps.
+To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and store it in temporal sequences.
 Subsequently, we extract the features from these 90-minute segments.
 In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values.
-Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (35; 40), prompting us to conduct separate analyses for these two categories.
+Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends [75, 72], prompting us to conduct separate analyses for these two categories.
 Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night.
@@
 For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources.
-This composite feature counts the number of notifications originating the Business, Communication, or Social app categories.
+This composite feature counts the number of notifications originating from the Business, Communication, or Social app categories.
 Notifications that do not belong to these app categories are counted as ‘app_noti’.
-5.2. Prediction Pipeline
-While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (58; 88).
+5.2 Prediction Pipeline
+While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds [76, 42].
 We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance.
@@
 In this project, we aim to develop regression models that accurately predict human emotions during home confinement.
-To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (72), Random Forest (RF) (71), and Gradient Boosting (GB) (21).
+To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) [77], Random Forest (RF) [78], and Gradient Boosting (GB) [79].
 KNN is a widely utilized model known for its versatility across various domains.
-It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable.
+It works by finding the K nearest neighbors to a given data point and using their values to predict the target variable.
 RF and GB, on the other hand, are ensemble models based on decision trees.
@@
 Feature Selection
-Various methods, such as SHAP (44), LIME (65), and the $F$-statistic ($F$) (60), can be used for feature selection.
+Various methods, such as SHAP [80], LIME [81], and the $F$-statistic ($F$) [82], can be used for feature selection.
 In this study, we employ the $F$-statistic to determine the appropriate number of features.
@@
 Validation
-To address the limited richness of our data, we implemented Nested Cross-Validation (CV) (50) to train and evaluate our prediction models.
-Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation (83) with different values of $k$.
+To address the limited richness of our data, we implemented Nested Cross-Validation (CV) [83] to train and evaluate our prediction models.
+Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation [84] with different values of $k$.
 The outer loop splits the data into $k$ folds, using each fold for testing.
@@
 Baselines and Metrics.
-Similar to previous human-centered studies (24; 82), we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value.
-To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (13).
+Similar to previous human-centered studies [44, 85], we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value.
+To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) [86].
 Generally, lower values of MAE or RMSE indicate better prediction performance.
-5.3. Prediction Result
-5.3.1. General Model
+5.3 Prediction Result
+5.3.1 General Model
 After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis were used to build the general models.
-Table 6.
-Prediction performance of the general models for valence and arousal with different feature sets.
+Table 6: Prediction performance of the general models for valence and arousal with different feature sets.
 The regression models are K-Nearest Neighbor (KNN), Random Forest (RF) and Gradient Boosting (GB), compared with Mean and Median baseline.
 The evaluation metrics are RMSE and MAE (with standard deviation).
-Prediction Result (RMSE(SD)/MAE(SD))
-Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
+RMSE/MAE(SD) Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
 All features
-[…]
-[…]
-Features with $F$-statistic ¿= 5
-[…]
-[…]
-Features with $F$-statistic ¿= 10
-[…]
-[…]
+Valence (29) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.963(.026)/0.776(.022) 0.903(.025)/0.717(.021) 0.925(.020)/0.743(.012)
+Arousal (29) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.088(.023)/0.844(.019) 1.005(.028)/0.766(.020) 1.084(.044)/0.756(.037)
+Features with $F$-statistic $\geq 5$
+Valence (17) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.956(.020)/0.769(.018) 0.904(.024)/0.717(.021) 0.986(.020)/0.725(.021)
+Arousal (14) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.064(.027)/0.818(.026) 1.015(.024)/0.778(.018) 1.087(.032)/0.747(.022)
+Features with $F$-statistic $\geq 10$
+Valence (13) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.936(.020)/0.730(.013) 0.912(.024)/0.725(.018) 0.987(.046)/0.724(.019)
+Arousal (12) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.064(.027)/0.819(.026) 1.011(.023)/0.774(.017) 1.076(.035)/0.745(.030)
 The prediction performance for the general models is presented in Table 6.
@@
 These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
-5.3.2. Group Models
-We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: $EARLY$ Workers (678 data points from 12 participants who typically commence work between 7:00 and 10:00 AM) and $LATE$ Workers (778 data points from 13 participants who begin work after 3:00 PM).
+5.3.2 Group Models
+We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: EARLY Workers (678 data instances from 12 participants who typically commence work between 7:00 and 10:00 AM) and LATE Workers (778 data instances from 13 participants who begin work after 3:00 PM).
 For the early workers, the average self-reported valence is 3.2 (±$\pm 0.9$) and arousal is 2.9 (±$\pm 1.0$).
@@
 (d) MAE for arousal
-Figure 7.
-Prediction performance of $valence$ and $arousal$ models for two user groups: $EARLY$ workers (typically starting work between 7:00 and 10:00 AM) and $LATE$ workers (starting work after 3:00 PM).
+Figure 7: Prediction performance of valence and arousal models for two user groups: EARLY workers (typically starting work between 7:00 and 10:00 AM) and LATE workers (starting work after 3:00 PM).
 The models are evaluated with 3-fold cross-validation, where the error bar presents the standard deviation.
@@
 Further insights will be explored in the later section (Section 5.4) by analyzing feature importance.
-5.3.3. Individual Models
+5.3.3 Individual Models
 The individual models for both valence and arousal were trained using all 29 features.
@@
 These findings contribute to our understanding of personalized prediction models for well-being, which could ultimately lead to more tailored interventions and support systems to enhance individuals’ well-being during periods of home confinement.
-(a) MAE for valence
-(b) RMSE for valence
-(c) MAE for arousal
-(d) RMSE for arousal
-Figure 8. Prediction performance of individual models using all features
-5.4. Feature Importance
+(a) valence (MAE)
+(b) valence (RMSE)
+(c) arousal (MAE)
+(d) arousal (RMSE)
+Figure 8: Prediction performance of individual models using all features
+5.4 Feature Importance
 To interpret the key factors influencing the prediction of valence and arousal during home confinement, we examine the feature importance of the best-performing models, which in this case are the RF regressors.
@@
 For brevity, we focus on the top 15 features for the general model and compare them with the top 5 features for the group models.
-Figure 9. Top 15 features based on impurity score with RF regressors of the general models.
+Figure 9: Top 15 features based on impurity score with RF regressors of the general models.
 As depicted in Figure 9, we observe that the most important features for both valence and arousal are similar, but with different weights.
@@
 Additionally, the day of the week (‘weekday’) impacts both models, although it has less importance in the arousal model.
-Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group models.
-It worth mentioning that ‘time_period’ is the most important feature for predicting $valence$ for $EARLY$ workers, but not for $LATE$ workers.
-These findings differ slightly from the observations reported in (35), which found that app usage patterns varied greatly by time and day.
+Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group model.
+It is worth mentioning that ‘time_period’ is the most important feature for predicting valence for EARLY workers, but not for LATE workers.
+These findings differ slightly from the observations reported in [75], which found that app usage patterns varied greatly by time and day.
 The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends.
-This aligns with the findings presented in (12), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
-Figure 10. Top 5 features based on impurity score with RF regressors of group models.
+This aligns with the findings presented in [87], which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
+Figure 10: Top 5 features based on impurity score with RF regressors of group models.
 Besides, as shown in Figure 10, for LATE workers, the same five features with similar importance distributions are observed for both valence and arousal predictions, all are related to app interaction.
-In contrast, $EARLY$ workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contribute to $valence$ prediction.
+In contrast, EARLY workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contributes to valence prediction.
 This difference may reflect the broader range of daytime activities available to EARLY workers, potentially explaining their lower prediction performance discussed in Section 5.3.2.
 For LATE workers, the reliance on app interaction highlights a potential concern regarding phone dependency.
-6. Discussion and Implications
-This study explores the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors.
+6 Discussion and Implications
+This study explores the impact of home confinement on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors.
 Although the study specifically focuses on the COVID-19 lockdown, we argue that its findings have broader relevance, especially in light of the growing trend of remote work and other situations of enforced isolation.
-6.1. Well-being and Social Roles During Home Confinement (RQ1)
+6.1 Well-being and Social Roles During Home Confinement (RQ1)
 Our analysis revealed that participants generally experienced a state of calmness and relaxation, reflected in relatively high levels of valence (positive emotional state) and lower levels of arousal (indicating a lack of intense emotional stimulation).
@@
 The segmentation of app usage data also identified key behaviors, such as increased use of communication apps and idle device time, which may reflect both increased virtual socialization and available free time.
-Although the findings are rooted in the context of the COVID-19 lockdown, the implications extend beyond this unique situation.
-Home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction.
+In wider contexts, home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction.
 The growing trend of telework and flexible work arrangements in the post-pandemic world further underscores the importance of understanding the emotional and behavioral impacts of prolonged confinement in various forms.
-6.2. Predicting Emotional States Through Mobile Usage (RQ2)
+6.2 Predicting Emotional States Through Mobile Usage (RQ2)
 In response to RQ2, we developed regression models that predict participants’ valence and arousal based on their mobile usage data.
@@
 Whether it’s remote workers, individuals recovering from surgery, or people experiencing limited mobility due to health conditions, the need for individualized and context-aware models remains a central takeaway.
-6.3. Design Implications
-Based on the findings of this study, we propose several design implications for future interventions and mobile applications aimed at supporting well-being during periods of home confinement or isolation:
-Personalized Predictive Models.
-The success of time-related features in predicting emotional states suggests that mobile apps should incorporate personalized, time-sensitive data to offer real-time emotional tracking and tailored interventions.
-For example, apps could adjust their notifications and suggestions based on users’ routines, ensuring that interventions are timely and relevant.
-Balanced Digital Interactions.
-Our study found that interruptions, particularly from communication apps, may exacerbate the blurring of work-life boundaries.
-Future apps should focus on helping users manage interruptions, offering customizable notification settings or “focus mode” features to reduce disruptions and help maintain better separation between work and personal life.
-Encouraging Healthy Digital Habits.
-Tools that support productivity, such as task managers and calendars, were associated with higher valence.
-Future designs should encourage healthy digital habits by integrating features that promote structure, goal-setting, and task management.
-Additionally, incorporating reminders for physical activity or mental health breaks could further support users’ emotional well-being during periods of isolation.
-Context-Aware Interventions.
-Given the differences in how users experience isolation based on their routines, digital interventions should be context-aware and adaptive.
-For example, users with later work hours may benefit from interventions encouraging physical activity or relaxation earlier in the day, while users with earlier start times may need more focused strategies for managing digital distractions during the day.
-6.4. Generalizability of the Findings
+6.3 Generalizability of the Findings
 Although this study specifically focuses on the COVID-19 home confinement context, we believe the findings can extend to a range of other isolation scenarios.
-While the emotional triggers—such as fear of infection in a pandemic or frustration with limited mobility in recovery—may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar.
+While the emotional triggers – such as fear of infection in a pandemic or frustration with limited mobility in recovery – may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar.
 Our study observed patterns like emotional detachment, stress, and blurred work-life boundaries, all of which are commonly experienced during prolonged isolation, regardless of the specific circumstances.
-The emotional dynamics of isolation—such as changes in social interactions, work-related stress, and adjusting to new routines—are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement.
+The emotional dynamics of isolation – such as changes in social interactions, work-related stress, and adjusting to new routines – are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement.
 For example, both remote workers and those recovering from surgery face challenges with work-life balance and social isolation.
 Our findings about app usage—especially productivity apps promoting positive emotional states—are relevant in these contexts, suggesting that digital tools can support well-being across various isolation scenarios.
-Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond COVID-19.
+Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond the studied context.
 While the types of apps and behaviors may vary, the principle of using mobile devices to track emotional states in real-time is adaptable to different situations, such as remote work or recovery periods.
@@
 These strategies can be adapted for any form of isolation, from remote work to recovery, offering broader applicability for improving mental health.
-7. Limitations
-This study has several limitations that must be acknowledged, which may affect the validity of our findings (78):
+7 Limitations
+This study has several limitations that must be acknowledged, which may affect the validity of our findings [88]:
 Time and Geographical Constraints (Threats to Construct and External Validity).
@@
 Device Usage and Untracked Data (Threats to Internal and Construct Validity).
-A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops while only contributing smartphone data.
+A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops, while only contributing smartphone data.
 These unmonitored interactions could have influenced emotional states and introduced noise into the data.
@@
 However, despite these challenges, the data collected was sufficient for analysis.
-8. Conclusion
-In this field study, we explored how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period.
-This rare event offers optimal conditions to while minimizing confounding factors like nature and physical social interactions.
+8 Conclusion
+In this field study, we explored how human emotions, social roles, and mobile usage behaviors were affected during the home confinement period.
+This rare event offers optimal conditions while minimizing confounding factors like nature and physical social interactions.
 It contributes to understanding the mental health implications of heavy reliance on technology and work-from-home arrangements.
 By tracking mental well-being, the study highlights how individuals manage work-life balance and adapt their schedules or arrangements (e.g., work-from-home or half-office half-home) effectively.
-Additionally, it provides valuable insights into psychological states and smartphone usage behaviours during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation.
+Additionally, it provides valuable insights into psychological states and smartphone usage behaviors during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation.
 Understanding how people navigate home confinement and the role of technology in their lives can guide the design of smart environments and remote support programs, improving care quality.
-These findings also have broader implications for understanding human behaviour during crises, supporting the development of policies and interventions for future challenges.
-Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviours during home confinement is crucial.
+These findings also have broader implications for understanding human behavior during crises, supporting the development of policies and interventions for future challenges.
+Despite the challenges we faced in collecting data during the lockdowns, we believe that gaining insights into human well-being and mobile usage behaviors during home confinement is crucial.
 This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health.
-Moving forward, our research will expand to investigate users’ desktop usage behaviour.
-By comparing the differences in usage behaviour across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
-9. Acknowledgments
+Moving forward, our research will expand to investigate users’ desktop usage behavior.
+By comparing the differences in usage behavior across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
+9 Acknowledgments
 We sincerely thank the anonymous reviewers. We state that this manuscript has no relation to our prior publications (or concurrently submitted papers).